A Complete Bibliography of ACM Transactions on Accessible Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 August 2024 Version 1.48

Title word cross-reference

3 [BLR^{+}15, BCH^{+}16, LRF^{+}20, MML21, SFSR^{+}22].

1 [Big17a, CRPA15]. ’14 [Ric16a]. 19 [FSR^{+}23, ZMN^{+}22].

2 [Big17b, LL15b, Ric16b, RCPA15]. 2009 [McC10]. 2011 [Yes12].

3 [PCRA15].

A3 [HKH^{+}09]. A3C [LKRV24]. AAC [FORB^{+}09, MW09]. AAC-Supported [FORB^{+}09]. Abilities [ASZ^{+}21]. Ability [NWVJ22, WKG^{+}11]. Ability-Based [NWVJ22, WKG^{+}11]. Abstract [KZM^{+}20]. Abuse [VRL^{+}23]. Academia [CLH24]. Academic [MTC22]. Acceptance [JW11, KHE^{+}17]. Access [BHKW22, Lad08, LA22, LCKG23, LHB11]. AccessComics2 [LJYO23].

Accessibility [AED^{+}22, CLH24, CM17, DH16, EŠR24, FSR^{+}23, FHG21, HAC^{+}15, KSV21, KLB14, MPPS23, MOL22, PDR^{+}16, PRM23, RZFW20, STK^{+}10, SW16, SBPW18, SJPW20, SST20, TFA17, TLHC09, Van08, YMMS22, YR19, ZZUK17, BYH12, SH12]. Accessible [BRDP21, GF17, HL10, KVH21, LJYO23, MMM^{+}20, MTC22, New08, PRR^{+}22, RBKC24, SJPW20, VAR^{+}19]. Accessing [ICL^{+}23, TCR^{+}15]. Accuracy [KKR^{+}23, MLG^{+}15, PKGG23]. Accurate [HL10]. ACE [TFA17]. ACM [BOK^{+}22]. Acquire [AWBV21]. Acquired [JLC^{+}22]. Acquisition [ABLM23]. across [GGK^{+}20]. Act [CLG22]. Acting [WGMM09]. Activity [FPC^{+}16, GRVE20, VGV^{+}21, VGVG23]. Adaptations [LCKG23]. Adapted [VIH13]. Adapting [WAPD08]. Adaptive [BWS^{+}19, BOK^{+}22, GPB09]. Addressing [MOL22]. Adult [LFM^{+}10, Sen16]. Adults [BGS^{+}16, GRVE20, GAY^{+}23, KDLD21, KZM^{+}20, LTH^{+}12, MAB^{+}24, MM09, SW09, VAR^{+}19, VSH^{+}21, VGV^{+}21, VGVG23, VRL^{+}23, WFL24]. Aesthetics [TFA17]. Affecting [KSV21]. Against [VRL^{+}23]. Age [CH16, DAL22, GGK^{+}20, SMB^{+}19, ORF12]. Age-friendly [SMB^{+}19]. Age-Related [DAL22]. Agency [TGMBS24]. Agent [WAMG22]. Agent-based [WAMG22]. Aging [CGH09, RWBM15, SM19]. AI [BK21, BCH^{+}21]. Aid [vEPM20]. Aids [HR22, RCPA15, SFW^{+}24]. AIGuide [LAL^{+}22]. Algorithms [SMB22]. Ally [LYB^{+}22]. ALT [MK22, EGBB23]. Alternative [PK10]. Alzheimer [FORB^{+}09, RWBM15]. Ambient [VCP^{+}15]. American [BHP19, HZGA08, Hue09, HGP^{+}17, KLH13, LH11]. Among [AED^{+}22, MS17, GRVE20]. Analysis [CH16, HAC^{+}15, KHE^{+}17, LGG^{+}20, MLG^{+}15, RZFW20, STK^{+}10, SHN^{+}11]. Analyzing [ZZS^{+}18]. and/or [YMMS22]. Android [RZFW20]. Animation [KLH13, KHE^{+}17, LH11]. Animations [Hue09]. Annotated [FM18]. Annotation [HKH^{+}09]. Anomalies [LFM15]. Aphasia [AMP08, GWM^{+}17, AM15, TNRW22]. Aphasia-Friendly [AM15]. App [ABLM23, JLC^{+}22, RZFW20, TMG23, VRL^{+}23]. Appeal [VGVG23]. Application [AMP08, vEPM20]. Applications [LFM^{+}10, LZL16, PK10, SVA^{+}18]. Applying [PMJB^{+}22]. Appreciation [PDR^{+}16]. Approach [ABWV22, BHH^{+}17, BBS^{+}21, NWVJ22]. Apps [CM17, KDLD21, YR19]. Apraxia [HBM^{+}21, PKL^{+}15]. AR/VR [BOK^{+}22]. Arithmetic [CKA18]. Armrests [CCC^{+}16]. Arousal [BDKK17]. Art [MPPS23, RCW^{+}18]. article [HS12]. Articulation [ATA15]. ASD [BWS^{+}19]. ASL [HZGA08]. Aspirations [SAL^{+}20]. Assess [HKH^{+}09]. Assessing [PMJB^{+}22, RJ10, WCG^{+}19]. Assessment [MLG^{+}15, PFM^{+}15, PS09]. Assessments [SMB^{+}19]. ASSETS [McC10, Ric16a, Yes12]. ASSETS’10 [HS11]. ASSETS’13 [LL15a, LL15b]. ASSETS’14 [Ric16b]. ASSETS’15 [Big17a, Big17b]. ASSETS’16 [HM18a]. ASSETS’17 [Bro19, Hue19]. ASSETS’18 [Mof20]. ASSETS’19 [SH21]. ASSETS’20 [KN22]. ASSETS’21 [Gue23]. Assist [SDO^{+}16]. Assistance [AMB^{+}20, AED^{+}22, LKE^{+}22, PHWH14, SSD^{+}22]. Assistant [SOG^{+}19]. Assisted [VCP^{+}15]. Assisting [PRR^{+}22]. Assistive [AAKS22, CRPA15, FRCH17, KMG^{+}18, MM24, MS17, PCRA15, RJ10, RCPA15, RWBM15, SFSR^{+}22, SW16, Van08, WAPD08, ZZUK17]. Association [LKRV24]. At-Home [PKGG23]. Ataxia [PKGG23]. Ataxia-Telangiectasia [PKGG23]. ATs [HKO^{+}23]. Attitudes [SMB^{+}19]. Audio [ASS22, BRDP21, CGR^{+}23, Göt18, HVKF20, KSV21, KDN23, LGG^{+}20, OBFK15, RBKC24]. Audio-Based [OBFK15]. Audio-enhanced [BRDP21]. Audio-Tactile [Göt18]. Audio-visual [KDN23]. Auditory [ASS22, JW11, MTM^{+}20, SDO^{+}16, TBC^{+}16, WM10]. Augment [MML21]. Augmentative [PK10]. Augmented [Göt18, LAL^{+}22, PMJB^{+}22, ZSS^{+}21]. Authentication [LKRV24, MFKL13]. Author [MTC22]. Authorable [TCR^{+}15]. Authoring [PWH11, TMG23]. Autism [BRH^{+}15, BBR^{+}17, FPC^{+}16, BBK^{+}18, MML21, PHT^{+}19, WSS^{+}17, ZFS^{+}18, ZZS^{+}18, ZAM^{+}21, ZSS^{+}21]. Autistic [AIB^{+}22, MAB^{+}24]. Automated [AB15, PHB23]. Automatic [KKR^{+}23, LFM15, MPPS23, PFM^{+}15, PHWH14, PSG22, SFW^{+}24]. Autonomous [BHP^{+}20, FHG21]. Avatar [AB15]. Avoidance [PAD^{+}21]. Aware [GCM15, VCP^{+}15]. Awareness [PDR^{+}16].

Balancing [CLG22, TFA17]. Barriers [KDLD21, MOL22, WFL24]. Based [AAKS22, APKE11, BLR^{+}15, DTD^{+}22, FPC^{+}16, GMM^{+}15, HKH^{+}09, HBM^{+}21, KMG^{+}18, KDN23, LFM15, LKRV24, MM09, NWVJ22, OBFK15, PKGG23, PK10, PS09, SVA^{+}18, WKG^{+}11, ZFS10, APS20, BBS^{+}21, BWS^{+}19, GTU^{+}20, WAMG22]. Basis [PSG22]. Battery [TRLW15]. Be [Sen16, PRM23]. Behavior [HKH^{+}09, MFKL13]. Behavioral [LHB11]. Behaviors [ZZS^{+}18]. Being [WGMM09]. Belt [vEPM20]. Best [PDR^{+}16, WDEA^{+}21]. Better [LYB^{+}22, PRM23]. between [HVKF20, IKTF22, MFKL13]. Beyond [HSKP22]. Bimanual [YJGF^{+}22]. Bit [TRLW15]. Blind [ASR^{+}20, CGR^{+}23, FM18, GGK^{+}20, GG16, HAC^{+}15, Her20, HR22, HVKF20, HSKP22, LJYO23, LYB^{+}22, LCKG23, MP13, NMG^{+}17, OF15, RVL^{+}17, SSG10, SOG^{+}19, SDO^{+}16, ZDW17, vEPM20]. Blindness [AMB^{+}20]. Book [LJYO23]. Brain [JLC^{+}22]. Broad [JLC^{+}22]. Broad-shallow [JLC^{+}22]. Browsing [KDN23, ZZUK17]. Brushing [PHWH14]. Build [LYB^{+}22]. Building [CLH24, PK10]. Bus [HAC^{+}15]. Business [MM24]. Bystanders [AAKS22].

C [ZAM^{+}21]. C-Hg [ZAM^{+}21]. Calibration [HL10]. Calls [VAR^{+}19]. Camera [AAKS22]. Camera-Based [AAKS22]. Cameras [SDO^{+}16]. Can [DAL22, EGBB23]. Cane [HSKP22]. Capabilities [PS09]. Capability [PFM^{+}15]. Capacity [CLH24, vEPM20]. Caption [MK22]. Captioning [MK22]. Captions [KH19, KLB14]. Capture [AB15, HL10]. Cards [SJPW20]. Care [SFSR^{+}22, WH14]. Career [AIB^{+}22]. Caregivers [HKO^{+}23]. Case [APKE11]. Catcher [ETK18]. Categorized [ATA15]. Cerebral [WBO^{+}09]. Challenges [EŠR24, FR15, FLKO10, MEL^{+}22, PDWT15]. Changes [AIP^{+}18, DAL22]. Characteristics [CH16]. Charts [DPS^{+}10]. CheerBrush [ZSS^{+}21]. Chief [Edi14, GL20, GL21, HM19]. Child [HBM^{+}21]. Childhood [PKL^{+}15]. Children [ABLM23, BDKK17, BRH^{+}15, CKA18, FPC^{+}16, FLKO10, HKO^{+}23, LCD17, PKGG23, PHT^{+}19, SSG10, SAL^{+}20, WBO^{+}09, ZZS^{+}18, ZAM^{+}21, ZSS^{+}21]. Chronic [YMMS22]. CIRVR [AIB^{+}22]. Classification [AIP^{+}18]. Classroom [EŠR24, KLB14, TBC^{+}16]. Classrooms [PDR^{+}16]. Clinical [BDKK17]. Clinician [DTD^{+}22]. Co [BBS^{+}21]. Co-designing [BBS^{+}21]. Coaching [ZSS^{+}21]. Coding [HKH^{+}09]. Cognitive [CH16, DAL22, HKO^{+}23, HF15, PHWH14, PHB23]. Collaborative [BRH^{+}15, MK22, STK^{+}10, WSS^{+}17, ZFS^{+}18, ZAM^{+}21, ZAM^{+}22]. Collection [HL10]. Color [FG12]. Colour [TFA17]. Combining [RNL17]. Comic [LJYO23, TNRW22, TNRW22]. Commentary [Edw08, Lad08, New08, Van08]. Common [BHKW22, Van08]. Communication [AMP08, BDKK17, GPB09, PK10, WAPD08]. Community [DM23]. CoMove [ZFS^{+}18]. Comparative [LBH23]. Comparing [SMB22]. Comparison [KZM^{+}20, MFKL13, OF15, WM10]. Compensation [AMB^{+}20]. Competency [BBS^{+}21]. Competency-based [BBS^{+}21]. Completion [SLHF10]. Complex [SHN^{+}11]. Component [PK10, TCR^{+}15]. Component-Based [PK10]. Computer [APS20, CB08, FLKO10, HKH^{+}09, HHW13, PKGG23, RNL17, SMB^{+}19, VIH13]. Computer-Based [HKH^{+}09, PKGG23]. Computers [Edw08, GY08, Lad08, New08, Van08]. Computing [AED^{+}22, BHH^{+}17, LKRV24, New08]. Concept [WKG^{+}11]. Conception [GPB09]. Concepts [MP13]. Conceptual [NWVJ22]. Conceptualizing [SW16]. Concerns [AAKS22, BHP^{+}20, MM24, SSD^{+}22, ZMN^{+}22]. Concurrent [GG16]. Conditions [YMMS22]. Conductance [BDKK17]. Conferencing [LBH23]. Confident [SW16]. conformance [BYH12]. Conscious [SW16]. Considerations [BDKK17]. Consistency [PBW17]. Content [GG16, KDN23, LHB11, MSDGO19, MK22, PRR^{+}22]. Context [CLH24, VCP^{+}15, WAPD08, ZMN^{+}22]. Context-Aware [VCP^{+}15]. Continuous [LZL16]. Controlled [RCW^{+}18, SMB22]. Convenient [LA22]. Conversation [ICL^{+}23]. Conversational [ICL^{+}23]. Conversations [FORB^{+}09]. Conversion [ATA15]. Cookie [CMT24]. Core [Van08]. cost [CKA18]. Could [BK21, Sen16, PRM23]. Covariance [WCH^{+}16]. COVID [FSR^{+}23, ZMN^{+}22]. COVID-19 [FSR^{+}23, ZMN^{+}22]. Crafting [BRDP21]. Created [TMG23]. Creating [ABWV22, FT18, ICL^{+}23, MMM^{+}20, MTC22]. Creation [EGBB23, TNRW22]. Creators [PRR^{+}22]. Criteria [MPPS23]. Critical [GRVE20]. Crossing [WG08]. Crossings [AMCM17]. Crosswalk [AMCM17]. Crowdsourcing [HAC^{+}15]. Cultivating [PDR^{+}16]. Cultural [LRF^{+}20]. Current [KDLD21, YR19]. Customizable [DTD^{+}22, LA22, ZSSA20]. Cyborg [EŠR24].

D [BLR^{+}15, BCH^{+}16, LRF^{+}20, MML21, SFSR^{+}22]. Daily [CRPA15, PCRA15, RCPA15]. Data [BK21, BLR^{+}15, FSR^{+}23, HL10, KMG^{+}18, LA22, LH11, VGVG23, WFL24]. Data-Driven [LH11]. Dataset [FM18, MAB^{+}24]. Datasets [BCH^{+}21]. Deaf [AED^{+}22, KH19, KDH21]. Deep [AMB^{+}20, ASC^{+}21, JLC^{+}22]. Dementia [DM23, WDEA^{+}21]. Demographic [KHE^{+}17]. Deploying [RBKC24]. Derived [KYO11, MST^{+}15]. Describing [HSKP22]. Description [CGR^{+}23]. Descriptions [HR22, MSM^{+}15]. Design [ASR^{+}20, ABWV22, ASS22, BHP19, BWS^{+}19, BRDP21, GTU^{+}20, HR22, LRF^{+}20, LBH23, LCD17, LCKG23, AM15, MPPS23, MTM^{+}20, NWVJ22, PDWT15, PHB23, PAD^{+}21, PSG22, RVL^{+}17, Sen16, SHN^{+}11, SBPW18, SJPW20, TFA17, VSH^{+}21, VRL^{+}23, WM10, WG08, WKG^{+}11, YJGF^{+}22, ZFS^{+}18, ZZS^{+}18]. Designed [AMP08, PHT^{+}19]. Designing [CLG22, DTD^{+}22, GMM^{+}15, GAY^{+}23, ICL^{+}23, KDLD21, PG17, SVA^{+}18, ZSSA20, BBS^{+}21]. Designs [KDH21]. Desires [PHT^{+}19]. Desktop [BHH^{+}17]. Detection [LFM15, MAB^{+}24, MSDGO19]. Development [AB15, KVH21, KVFO^{+}23, LBH23, LCD17, PKL^{+}15, PRR^{+}22, PS09, WAMG22]. Developmental [SVA^{+}18, VRL^{+}23]. Device [GPB09, LKRV24]. Devices [LTH^{+}12, PG17, SC20, SFSR^{+}22]. Devising [CKA18]. Dialog [WGMM09]. Diary [SW16]. Dictionary [HAGH21]. Differences [HVKF20, ORF12]. Differentiation [FG12]. Digital [AMP08, CLH24, FL23, GG16, PWH11]. Digital-Paper [PWH11]. Dimensional [DPS^{+}10]. Directional [SDO^{+}16]. Directions [FHG21, WG08]. Disabilities [ASZ^{+}21, BSB22, BK21, Edw08, GY08, HKO^{+}23, HF15, Lad08, LCD17, New08, PHWH14, PHB23, SVA^{+}18, Van08, VRL^{+}23, YMMS22, ZMN^{+}22]. Disability [BBS^{+}21, FRCH17, ZFS10]. Disabled [SBPW18, KDN23]. Disease [FR15, FORB^{+}09, RWBM15, TST^{+}24]. Disorder [BRH^{+}15, BBR^{+}17, ZFS^{+}18, ZAM^{+}21, ZSS^{+}21]. Disordered [PFM^{+}15]. Disorders [ATA15, WSS^{+}17]. Displaying [KLH13]. Displays [LKE^{+}22]. Distance [SMB22]. Distinguishing [HHMT13]. Diverse [MEL^{+}22, PDR^{+}16]. DIY [HKO^{+}23]. DIY-ATs [HKO^{+}23]. Do [CLG22]. Doesn’t [WGMM09]. Don’t [MM24]. Down [AIP^{+}18, WFL24, FLKO10, KDLD21, MFKL13]. Drain [TRLW15]. Drawing [ASR^{+}20]. Driven [LH11]. Driving [BWS^{+}19]. Drop [AIP^{+}18]. Drop-Down [AIP^{+}18]. During [AMB^{+}20, FSR^{+}23, KLH13, MAB^{+}24]. Dwell [PDWT15]. Dwell-Free [PDWT15]. Dyadic [WSS^{+}17]. Dysarthric [LFM15, MLG^{+}15]. Dysfunction [WAMG22]. Dyslexia [RBY16].

E-Mail [BGS^{+}16]. Easy [CKA18]. Easy-to-use [CKA18]. Echolocation [AWBV21, ABWV22]. Echolocation-Enabled [ABWV22]. Economically [CKA18]. Ecosystems [SFSR^{+}22]. Editorial [BF09, GL21, HM19, McC10, SH13, Tre08]. Editors [Edi14, GL20, GL21, HM19]. Editors-in-Chief [Edi14, GL20, GL21, HM19].

Education [BCH^{+}16, CLH24, KVH21, PRM23]. Educational [FL23, KDH21, MM24]. Effect [DAL22, FORB^{+}09, HAGH21, KLH13, RBY16, TBC^{+}16]. Effective [PMJB^{+}22, Sen16]. Effects [HHW13, BBK^{+}18, LJYO23, PMJB^{+}22, SMB^{+}19, SM19, TMY^{+}09]. Efficient [MSDGO19]. Electronic [CLG22, HR22, vEPM20]. elusive [BYH12]. Email [AM15]. Emergency [VAR^{+}19]. Emotional [BDKK17, WNI^{+}18]. Empirically [VSH^{+}21]. Employers [AIB^{+}22]. Empower [VRL^{+}23]. Empowerment [Lad08]. Enabled [ABWV22, SC20]. Enabling [TNRW22]. Encoding [KSV21]. Engaging [MP13, SJPW20]. English [KH19]. enhanced [BRDP21]. Enhancement [ZSSA20]. Enhancing [ASZ^{+}21]. Enriching [DM23]. Enterprise [SST20]. Entry [SC20]. Environment [FPC^{+}16, TBC^{+}16, ZFS^{+}18]. Environmental [KSV21]. Environments [ABWV22, MTM^{+}20, MEL^{+}22]. Epidemiology [RZFW20]. Epidemiology-inspired [RZFW20]. Equality [BK21]. Errors [AMB^{+}20]. EVA [GWM^{+}17]. Evaluating [BRH^{+}15, KDH21, SDO^{+}16, TRLW15, WBO^{+}09, ZSSA20]. Evaluation [AB15, AMP08, ASS22, BHP19, CMT24, HBM^{+}21, HZGA08, HGP^{+}17, KLH13, KYO11, KLB14, MPPS23, MGBP17, MML21, PG17, PRR^{+}22, PS09, RVL^{+}17, SSB^{+}15, SLHF10, TMG23, VCP^{+}15, ZFS^{+}18]. Every [BHKW22]. Everyday [NMG^{+}17, ORF12]. Examples [WKG^{+}11]. Excel [DPS^{+}10]. Executive [WAMG22]. Exemplars [ATA15]. Exercise [RVL^{+}17]. Exergame [RVL^{+}17]. Expanding [BRL21]. Expansion [HHW13]. Experience [ETK18, ICL^{+}23, KHE^{+}17, BBK^{+}18, LJYO23, PRM23, ORF12]. Experienced [DAL22]. Experiences [AED^{+}22, FSR^{+}23, FR15, LBH23, PWC^{+}22]. Experiencing [GWM^{+}17]. Experimental [LGG^{+}20, TMG23]. Experimentation [GPB09]. Expert [CLH24]. Exploration [GRVE20, KDN23, MGBP17]. Exploratory [STK^{+}10]. Explore [RCW^{+}18]. Exploring [BGS^{+}16, BHP^{+}20, CCC^{+}16, IKTF22, KDLD21, LRF^{+}20, MM24, MM09, TGMBS24, TST^{+}24, TBC^{+}16, TLHC09, VGV^{+}21, VGVG23, MK22]. expression [TNRW22]. Extended [HAC^{+}15, SMB^{+}19]. Extensible [RBKC24]. Extraction [WNI^{+}18]. Extremity [LKRV24]. Eye [PDWT15, VIH13]. Eye-Tracking [VIH13]. Eyes [DH16, RVL^{+}17]. Eyes-Free [RVL^{+}17].

Face [RNL17]. Facilitating [BHKW22]. Factor [MLG^{+}15]. Factors [KHE^{+}17, KSV21]. Fairness [BK21]. Family [CCC^{+}16]. Fast [RNL17]. FATE [BCH^{+}21]. Feasibility [WNI^{+}18]. Features [BHKW22, SOG^{+}19]. Feedback [HGP^{+}17, OBFK15, TGMBS24]. Field [AMP08, LTH^{+}12, AM15]. Files [PRR^{+}22]. Filteryedping [PDWT15]. Find [WCG^{+}19]. Fine [ZZS^{+}18, ZAM^{+}21]. Finger [SDO^{+}16]. Finger-Mounted [SDO^{+}16]. Finland [SAL^{+}20]. Fit [RJ10]. Floor [MMM^{+}20]. Font [RBY16]. Formant [MST^{+}15]. Formant-Derived [MST^{+}15]. Format [PG17]. Frame [TRLW15]. Framework [CKA18, LKRV24, PK10, RBKC24]. Free [PDWT15, RVL^{+}17, VAR^{+}19]. Friendly [FT18, AM15, SMB^{+}19]. Fully [AB15, FHG21]. Functional [PS09]. Future [FLKO10, FHG21, LCKG23, New08].

Gadgets [VGV^{+}21]. Game [BRH^{+}15, DTD^{+}22, SSG10]. Games [APKE11, GMM^{+}15, LCD17, VIH13]. Gaze [RNL17, SMB22]. Gaze-Controlled [SMB22]. Generating [WBO^{+}09]. Generation [CGR^{+}23, HZGA08]. Geometrical [MP13]. Germany [FL23]. Gest [CCC^{+}16]. Gest-Rest [CCC^{+}16]. Gesture [DTD^{+}22, RCW^{+}18, SVA^{+}18]. Gesture-Based [DTD^{+}22, SVA^{+}18]. Gesture-Controlled [RCW^{+}18]. Gestures [OBFK15, RNL17]. Gets [EGBB23]. GIS [AMCM17]. Glasses [TBB20]. Global [SFSR^{+}22]. Glove [HL10]. Goal [WG08]. Google [HAC^{+}15]. Government [SHN^{+}11]. Graphics [BMD^{+}16, GTU^{+}20, Göt18, MML21, PG17, RBKC24]. Graphing [TBC^{+}16]. Graphs [TBC^{+}16, WM10]. Grassmann [WCH^{+}16]. Greetings [Edi14, GL20]. Grid [ZFS10]. Grid-Based [ZFS10]. Gripper [ZZS^{+}18, ZAM^{+}21, ZAM^{+}22]. Grounded [VSH^{+}21]. Group [MP13, TCR^{+}15]. Guest [BF09, McC10, Tre08]. Guidance [LAL^{+}22, SDO^{+}16]. Guideline [HKH^{+}09]. Guidelines [FT18, GAY^{+}23, GTU^{+}20, LRF^{+}20, LBH23, SVA^{+}18, VSH^{+}21]. Guiding [MSM^{+}15].

Hand [LAL^{+}22, OF15, TGMBS24]. Handheld [AS10, SC20]. Hands [VAR^{+}19, SGO17]. Hands-free [VAR^{+}19]. Haptic [AS10, BLR^{+}15, KYO11, MP13, PG17, RBKC24, SDO^{+}16, ZZS^{+}18, ZAM^{+}21, ZAM^{+}22]. Haptic-Gripper [ZZS^{+}18, ZAM^{+}21, ZAM^{+}22]. Hard [AED^{+}22, KH19, KDH21]. Hard-of-Hearing [AED^{+}22, KDH21]. HCI [GRVE20, HKH^{+}09]. Head [LKE^{+}22, ZSSA20]. Head-Mounted [LKE^{+}22, ZSSA20]. Health [SHN^{+}11, WFL24, ZMN^{+}22]. Hear [CLG22]. Hearing [AED^{+}22, KH19, KSV21, KDH21]. Helping [PHB23]. Hg [ZZS^{+}18, ZAM^{+}21, ZAM^{+}22]. Highlighting [KDH21]. Hindering [PHB23]. Histological [ZDW17]. Hockey [CLG22]. Home [PKGG23, RWBM15]. Homes [WH14]. House [VAR^{+}19]. HRI [SMB22]. HTML [AIP^{+}18, KYO11]. Human [BRL21, KLH13, RNL17, SLHF10]. Human-Computer [RNL17]. Humanizing [SBPW18].

ID [FPC^{+}16]. Ideas [LCKG23]. Identification [AIP^{+}18]. Identifying [LHB11, SGOM14]. I’m [YMMS22]. Image [AMP08, LKRV24, MSM^{+}15, RBKC24]. Image-Association-Based [LKRV24]. Imagery [AMCM17, EGBB23]. Images [ZDW17]. Immersive [DTD^{+}22, ETK18, VSH^{+}21, VIH13]. Impact [BRH^{+}15, WCG^{+}19, ZMN^{+}22, ZFS10]. Impacts [KVFO^{+}23]. Impaired [AAKS22, CGR^{+}23, CKA18, CB08, Göt18, LRF^{+}20, LKE^{+}22, LHB11, MMM^{+}20, PWC^{+}22, SGO17, WM10, ZDW17, ZZUK17]. Impairment [ABLM23, AWBV21, APS20, FHG21, KMG^{+}18, MTM^{+}20]. Impairments [AMB^{+}20, AS10, BHKW22, BDKK17, BLR^{+}15, BHP^{+}20, CMT24, GF17, GBK20, KSV21, LKRV24, LGG^{+}20, MS17, MOL22, MEL^{+}22, PKGG23, PAD^{+}21, SC20, TGMBS24, TBB20, WG08]. Imperfect [KH19]. Implementation [SSB^{+}15]. Implementing [NWVJ22]. Implications [BDKK17, BCH^{+}16, PRM23, SHN^{+}11]. Improve [APS20, ICL^{+}23, KVH21, LFM^{+}10, LHB11, MM09, SSG10]. Improved [HR22]. Improvement [STK^{+}10]. Improves [JW11]. Improving [CM17, HAC^{+}15, LCKG23, ZZUK17]. Improvisation [IKTF22]. In-Page [SFW^{+}24]. In-Person [DM23]. Inaudible [CMT24]. INC-Hg [ZAM^{+}22]. Inclusion [HKO^{+}23]. Inclusive [BOK^{+}22, BRL21, EGBB23, IKTF22, LBH23, PHB23]. Inclusivity [EŠR24]. Independent [GBK20, KDLD21]. Independently [MEL^{+}22]. Index [JW11]. India [SVA^{+}18, SAL^{+}20]. Individuality [ATA15]. Individuality-Preserving [ATA15]. Individuals [AIB^{+}22, BWS^{+}19, BBR^{+}17, DAL22, BBK^{+}18, MFKL13, RWBM15, SVA^{+}18, WSS^{+}17, WNI^{+}18]. Indoor [GGK^{+}20, MEL^{+}22, SOG^{+}19, TMG23]. Industry [PRM23]. Inertial [FM18]. Inference [WFL24]. Inference-making [WFL24]. Inflected [LH11]. Influencing [KHE^{+}17]. Information [CH16, CB08, CGH09, Her20, HF15, SST20, WNI^{+}18, WCG^{+}19]. Initial [LFM^{+}10]. Injury [JLC^{+}22]. Input [CCC^{+}16, HVKF20, OF15, SM19, YJGF^{+}22]. Inquiry [BRDP21]. Insights [ASR^{+}20, DM23, KMG^{+}18]. inspired [RZFW20]. Instructions [FT18]. Instruments [FL23]. Intellectual [ASZ^{+}21, BSB22, BBS^{+}21, VRL^{+}23]. Intelligent [ZAM^{+}22]. Intelligibility [MLG^{+}15, TRLW15]. Interact [WGMM09]. Interaction [APS20, CRPA15, LA22, PCRA15, RNL17, RCPA15, RWBM15, SLHF10, TCR^{+}15, TMY^{+}09, WSS^{+}17]. Interactions [YJGF^{+}22, ORF12]. Interactive [CM17, LRF^{+}20, PHT^{+}19, RCW^{+}18, ZSS^{+}21]. Interdisciplinary [BCH^{+}21]. Interests [AED^{+}22]. Interface [BGS^{+}16, ICL^{+}23, LGG^{+}20, VCP^{+}15, VAR^{+}19, VIH13]. Interfaces [ASR^{+}20, JLC^{+}22, LFM^{+}10]. Internet [ASZ^{+}21]. Intervention [HKH^{+}09, TCR^{+}15, WSS^{+}17, ZFS^{+}18]. Interventions [YMMS22]. Interview [AIB^{+}22]. Interviews [MAB^{+}24]. Introduction [Big17a, Big17b, Bro19, CRPA15, CGH09, GBK20, Gue23, HS11, HS12, HM18a, HM18b, Hue19, KN22, LL15a, LL15b, MW09, McG12, Mof20, PCRA15, RE24, Ric16a, Ric16b, RCPA15, SH08, SH21]. Investigating [BCH^{+}16, HF15, MFKL13, NMG^{+}17, WDEA^{+}21, YMMS22, ZFS10]. Investigation [DPS^{+}10]. Investigations [LTH^{+}12]. Invisible [CMT24]. Involving [SMB22]. iPad [BRH^{+}15]. Irrelevant [VGV^{+}21]. Isolated [WCH^{+}16]. Issue [Big17a, Big17b, BOK^{+}22, Bro19, CRPA15, CGH09, GBK20, Gue23, HS11, HM18a, HM18b, Hue19, KN22, LL15a, LL15b, MW09, McG12, Mof20, RE24, Ric16a, Ric16b, SH21, Yes12]. IssuePart [PCRA15]. Iterative [AM15, PAD^{+}21]. ITHACA [PK10].

Job [MAB^{+}24]. Journey [ASS22]. Just [YMMS22].

Keeping [Edw08]. Kenyan [HKO^{+}23]. Key [AB15, KDH21]. Keyboard [LYB^{+}22]. Kinesthetically [AS10]. Knowledge [AWBV21, HR22, ORF12].

Labelling [SFW^{+}24]. Laboratory [HHMT13, NMG^{+}17]. Landmark [HAC^{+}15]. Landscape [BCH^{+}21]. Language [ASC^{+}21, BCH^{+}21, CRPA15, HL10, HGP^{+}17, KHE^{+}17, LZL16, MSDGO19, PWH11, PCRA15, PSG22, RCPA15, SGOM14, TRLW15, WCH^{+}16, HAGH21, BHP19, HZGA08, Hue09, HGP^{+}17, KLH13, LH11]. Lanka [ASZ^{+}21]. Large [BRL21, KMG^{+}18, PG17, RZFW20, SOG^{+}19]. Large-Format [PG17]. Large-Scale [KMG^{+}18, RZFW20, SOG^{+}19]. Law [BK21]. Layered [LFM^{+}10]. Leading [ASR^{+}20, GRVE20]. Learn [LTH^{+}12]. Learnability [LFM^{+}10]. Learning [AMB^{+}20, ASC^{+}21, CKA18, LCKG23, MAB^{+}24, ZMN^{+}22]. Lecture [KDH21]. Less [CKA18]. Less-privileged [CKA18]. Lifecycle [MEUHB23]. Limb [ETK18]. Limited [YJGF^{+}22]. Line [IKTF22]. Linguistically [Hue09]. Listeners [WM10]. Listening [BRL21, IKTF22]. Literacy [ABLM23, WFL24]. Literature [MOL22]. Living [VCP^{+}15]. Localization [AMCM17]. Location [GCM15]. Location-Aware [GCM15]. Locations [HAC^{+}15, SAL^{+}20]. Locomotion [VIH13]. Longitudinal [HBM^{+}21, KMG^{+}18]. Look [WCG^{+}19]. Look-up [WCG^{+}19]. Loss [KSV21, PFM^{+}15]. Low [CKA18, LA22, LCKG23, RVL^{+}17, TRLW15, ZSSA20]. Low-cost [CKA18]. Low-Vision [RVL^{+}17, LA22].

Machine [CGR^{+}23, MAB^{+}24]. Machines [Sen16]. Made [EGBB23]. Magnifier [LA22]. Mail [BGS^{+}16]. Mainstream [FPC^{+}16]. Maintenance [SMB22]. Making [MSM^{+}15, PRR^{+}22, SSB^{+}15, VAR^{+}19, WFL24]. many [LCKG23]. Map [TMG23]. Mapping [YJGF^{+}22]. Mappings [KYO11, WM10]. Maps [Her20, TMG23]. Marker [BDKK17]. Materials [PWC^{+}22, PWH11]. Mathematical [KDN23]. Matrices [WCH^{+}16]. Mean [WGMM09]. Means [AWBV21]. Measurement [PS09]. Measures [WSS^{+}17]. Measuring [GCM15, KKR^{+}23]. Mechanism [SC20]. Medications [TST^{+}24]. Melodie [BRDP21]. Memory [vEPM20]. Mental [Her20, ZMN^{+}22]. Menu [AIP^{+}18, JW11, MM09]. Message [HM19]. Messaging [LYB^{+}22]. Metacognitive [WAMG22]. MetaCogs [WAMG22]. Method [SJPW20]. Methodological [BDKK17]. Methodology [APS20, KYO11, PS09]. Methods [ASC^{+}21, HGP^{+}17, MFKL13, MM09, PG17]. Mice [WG08]. Micro [Van08].

Micro-Assistive [Van08]. Microsoft [DPS^{+}10]. Mild [WDEA^{+}21]. Mind [AMCM17, RCW^{+}18]. Mining [AMCM17]. Mitigating [WAMG22]. Mixed [MTM^{+}20]. Mobile [ABLM23, AMP08, CM17, KDLD21, LFM^{+}10, LTH^{+}12, McG12, SC20, TRLW15, YR19, ZZUK17]. Mobility [GBK20, HSKP22, KMG^{+}18, MEL^{+}22, TBB20]. Modalities [CM17]. Model [Hue09, LRF^{+}20]. Modeling [LZL16, NWVJ22, SW09, WAMG22]. Models [FG12]. Moderate [WDEA^{+}21]. Modulation [MST^{+}15]. Motion [APKE11, AB15, HL10]. Motion-Based [APKE11]. Motion-Capture [HL10]. Motivated [Hue09]. Motor [APS20, LCD17, PKGG23, TGMBS24, TLHC09, WG08, ZZS^{+}18, ZAM^{+}21]. Mounted [LKE^{+}22, SDO^{+}16, ZSSA20]. Mouse [KZM^{+}20]. MouseClicker [TGMBS24]. Movement [GMM^{+}15, YJGF^{+}22]. much [PRM23]. Multi [GWM^{+}17, LFM^{+}10, TMG23]. Multi-Layered [LFM^{+}10]. Multi-scale [TMG23]. Multi-User [GWM^{+}17]. Multimodal [BHH^{+}17, CB08, DPS^{+}10, GTU^{+}20, PWH11, SSG10, ZDW17]. Multiple [PMJB^{+}22]. Music [FL23, IKTF22, LCKG23, MML21]. Musical [FL23].

N [PWH11]. Narrow [JLC^{+}22]. Narrow-deep [JLC^{+}22]. Native [HZGA08]. NavCog3 [SOG^{+}19]. Navigating [EŠR24, TLHC09]. Navigation [AMB^{+}20, GGK^{+}20, JW11, SSG10, SFW^{+}24, SOG^{+}19, WCG^{+}19, ZFS10]. Needs [BHP^{+}20, FL23, MEL^{+}22, SAL^{+}20, SST20]. Network [KSV21]. Networking [GAY^{+}23]. Neurodevelopmental [BDKK17]. Neuromuscular [FR15]. Neurotypical [MFKL13, WNI^{+}18]. Noise [KSV21]. Non [CLG22, MGBP17, SMB^{+}19]. Non-Visual [MGBP17]. Non-Visually [CLG22]. None [MM24]. Nontraditional [RJ10]. Nonverbal [HKH^{+}09]. Nonvisual [BHH^{+}17, OF15, PG17]. Notices [CMT24]. Novel [ETK18, KYO11, MGBP17, MAB^{+}24, ZSS^{+}21]. Novice [MSM^{+}15].

Obstacle [PAD^{+}21]. OCR [CM17]. Office [DPS^{+}10]. Old [WGMM09]. Older [BGS^{+}16, GRVE20, GAY^{+}23, HHW13, KZM^{+}20, LFM^{+}10, LTH^{+}12, McG12, MM09, Sen16, SHN^{+}11, SMB^{+}19, SW09, VAR^{+}19, VSH^{+}21, VGV^{+}21, VGVG23, WGMM09]. On-Hand [OF15]. On-Phone [OF15]. One [LCD17, YJGF^{+}22]. One-Switch [LCD17]. Online [DM23, MMM^{+}20, ZMN^{+}22]. Onscreen [LYB^{+}22]. Open [PK10, RBKC24, DPS^{+}10]. Open-Source [RBKC24]. Operating [FR15]. Opportunities [KVH21, KDLD21, MEUHB23]. Orientation [GBK20, KMG^{+}18, TBB20]. oriented [ASR^{+}20]. Orienting [AS10]. Outgoing [HM19]. Output [FORB^{+}09]. Overall [PCRA15]. Overcome [FR15]. Overviews [ASS22]. Overwhelmed [YMMS22].

Page [SFW^{+}24]. Palette [TFA17]. Palsy [WBO^{+}09]. Panning [PG17]. Paper [PWH11]. Papers [MTC22]. Parents [PHT^{+}19]. Park [GWM^{+}17].

Parkinson [TST^{+}24]. Part [Hue19, Big17a, Big17b, CRPA15, LL15b, Ric16b, RCPA15]. Partially [Her20, HR22]. Participants [WNI^{+}18]. Participatory [ABWV22]. Past [New08]. Patients [PMJB^{+}22]. Pattern [PBW17]. Patterns [WNI^{+}18]. Pausing [Hue09]. PC [FR15]. PDF [PRR^{+}22]. Peer [TCR^{+}15]. Pen [MM09]. Pen-Based [MM09]. People [AMB^{+}20, AMP08, AS10, AWBV21, ASZ^{+}21, APS20, BSB22, BHKW22, BBS^{+}21, BK21, BLR^{+}15, CGR^{+}23, Edw08, FHG21, GWM^{+}17, GY08, Göt18, GF17, GG16, GBK20, Her20, HR22, HSKP22, HF15, JLC^{+}22, KMG^{+}18, KH19, KSV21, Lad08, LJYO23, LRF^{+}20, LKRV24, LKE^{+}22, LGG^{+}20, LYB^{+}22, MML21, MOL22, MEL^{+}22, New08, PHWH14, PAD^{+}21, RBY16, SC20, SBPW18, SDO^{+}16, TGMBS24, TNRW22, TST^{+}24, TBB20, Van08, WG08, WDEA^{+}21, YMMS22, ZSSA20]. Perceive [GG16]. Perceived [VGVG23]. Perception [MPPS23, ZDW17]. Perceptions [AED^{+}22]. Performance [GCM15, GGK^{+}20, HAGH21, HHMT13, HHW13, JW11, NMG^{+}17, OF15, PDWT15, PS09, VGVG23, WG08]. Performance-Based [PS09]. Performing [VIH13]. Person [DM23]. Personal [LCKG23, RWBM15]. Personalized [ZZUK17]. Persons [BHP^{+}20, FORB^{+}09, MTM^{+}20, SHN^{+}11, ZDW17]. Perspective [BSB22, BCH^{+}21, FL23]. Perspectives [CRPA15, CLH24, DTD^{+}22, HKO^{+}23, PCRA15, RCPA15]. Phonated [MST^{+}15]. Phone [LFM15, OF15]. Phone-Based [LFM15]. Phoneme [ATA15]. Phoneme-Categorized [ATA15]. Phones [BGS^{+}16]. Physical [GRVE20, PS09, TGMBS24, VGVG23, YMMS22, ZFS10]. Physiological [MAB^{+}24]. Physiology [BWS^{+}19]. Physiology-based [BWS^{+}19]. Pictures [RCW^{+}18]. Pitch [MST^{+}15]. Planning [ASS22]. Plans [MMM^{+}20]. Platform [AIB^{+}22, BWS^{+}19]. Plausible [MST^{+}15]. Playback [CB08]. Playing [CLG22, IKTF22]. Pleasure [VGVG23]. Plugins [MEUHB23]. Pointing [HHMT13, PKGG23, RNL17]. Policy [FHG21]. Pool [DH16]. Possibilities [CCC^{+}16]. Practice [KVH21, PWC^{+}22, PRM23]. Practices [FSR^{+}23, PDR^{+}16, WDEA^{+}21]. Predicting [KH19]. Prediction [GCM15, MLG^{+}15, TMY^{+}09]. Predictors [MS17]. Preferences [BHP^{+}20]. Presentation [DPS^{+}10]. Preserving [ATA15]. Principles [PG17, WKG^{+}11]. Print [KDN23]. Print-disabled [KDN23]. Printed [SFSR^{+}22, SDO^{+}16]. Printing [BCH^{+}16]. prior [ORF12]. Priorities [MK22]. Privacy [AAKS22, MM24, SSD^{+}22]. privileged [CKA18]. Problem [SHN^{+}11]. Process [WSS^{+}17]. Processes [EGBB23]. Processing [PSG22]. Product [FT18]. Productivity [ASR^{+}20]. Productivity-oriented [ASR^{+}20]. Professionals [AED^{+}22]. Proficiency [MS17, SMB^{+}19]. Program [TCR^{+}15]. Programming [MOL22]. Programs [DM23]. Progress [KVFO^{+}23]. Project [ETK18]. Prompting [PHB23]. Proof [SLHF10]. Properties [BBK^{+}18]. property [BYH12]. Prosthetic [LAL^{+}22]. Protection [BK21]. Prototyping [VRL^{+}23]. Psychometric [BHP19]. Public [HAC^{+}15]. Puck [CLG22]. Pun [WBO^{+}09]. Pupils [MP13]. Puzzle [BHKW22]. PVI [AWBV21].

Qualitative [VCP^{+}15]. Quality [TMY^{+}09]. Quantification [PKGG23]. Quantitative [VCP^{+}15]. Questionnaires [BHP19]. Quo [FL23].

Range [DAL22]. Rate [MLG^{+}15]. Rates [BRL21, TRLW15]. Readability [RBY16]. Reader [FSR^{+}23, KVH21, KVFO^{+}23, LJYO23, LBH23, MEUHB23, SFW^{+}24, WCG^{+}19]. Readers [KDN23, MMM^{+}20]. Readiness [AIB^{+}22]. Reading [AED^{+}22, RCPA15, SDO^{+}16]. Reading-Assistance [AED^{+}22]. Real [HHMT13, KZM^{+}20, LZL16, MML21, RVL^{+}17]. Real-time [MML21]. Real-World [HHMT13, KZM^{+}20, LZL16, RVL^{+}17]. Reality [AIB^{+}22, BWS^{+}19, BBR^{+}17, DTD^{+}22, ETK18, BBK^{+}18, LAL^{+}22, MTM^{+}20, PMJB^{+}22, TBB20, VSH^{+}21, ZZS^{+}18, ZAM^{+}22, ZSS^{+}21]. Recognition [KKR^{+}23, LZL16, SGO17, WCH^{+}16, HAGH21]. Recognizer [MLG^{+}15]. Recommendations [GTU^{+}20]. Reconstruction [MST^{+}15]. Records [LA22]. Reflecting [SJPW20]. Reflections [GRVE20, MTC22]. Refreshable [PG17]. Regarding [BHP^{+}20]. Regression [KHE^{+}17]. Rehabilitation [BBR^{+}17, DTD^{+}22, ETK18, PMJB^{+}22]. Related [DAL22]. Relationships [BRH^{+}15]. Reliability [PKGG23, BYH12]. Reliefs [RCW^{+}18]. Reminding [JLC^{+}22]. Remote [PKL^{+}15, PMJB^{+}22, WDEA^{+}21]. Renderings [RBKC24]. Reports [PHT^{+}19]. Representation [BLR^{+}15]. Representations [HR22]. Representative [DH16]. Representing [SH12]. Requirements [FL23]. Research [DH16, FLKO10, HKH^{+}09, SH12]. Residential [WH14]. Rest [CCC^{+}16]. Retrieval [CH16]. Review [GAY^{+}23, LKE^{+}22, MOL22, PSG22]. Reviewers [BAS^{+}09, oACS11]. Reviewing [HVKF20]. Riders [HAC^{+}15]. Road [TBB20]. Robots [EŠR24, RWBM15]. Roles [EGBB23]. Rotation [AMB^{+}20]. Route [ASS22, HR22]. Rude [MM24].

Safe [SMB22]. Scalable [LZL16]. Scale [KMG^{+}18, RZFW20, SOG^{+}19, TMG23]. Scanning [GPB09, GG16, WNI^{+}18]. Scenarios [SJPW20]. Scenes [HSKP22]. Schedule [FPC^{+}16]. Schizophrenia [WNI^{+}18]. Schools [FL23]. Sclerosis [PMJB^{+}22]. Scoping [LKE^{+}22]. Screen [FSR^{+}23, KVH21, KVFO^{+}23, LA22, LBH23, MEUHB23, RBY16, SFW^{+}24, WCG^{+}19]. Screen-Reader [SFW^{+}24]. Search [ASZ^{+}21, HAGH21, HF15]. Security [MM24]. Seeing [DH16]. Seek [WCG^{+}19]. Selection [HHW13, MM09, SM19]. Self [SW16, TNRW22, TST^{+}24]. Self-Confident [SW16]. Self-Conscious [SW16]. Self-expression [TNRW22]. Self-track [TST^{+}24]. Semantic [SOG^{+}19]. Senior [FT18]. Senior-Friendly [FT18]. Sensor [FM18]. Sensory [DAL22, Her20, KDN23]. Series [FM18]. Sessions [MML21]. Set [PBW17]. Settings [MM24]. Severe [LCD17]. shallow [JLC^{+}22]. Shape [BK21, CB08, MGBP17]. Shared [AAKS22]. Sharing [DM23, MSDGO19, SGOM14]. Sibylle [WAPD08]. Sighted [ASR^{+}20, AAKS22, GGK^{+}20, GG16, Her20, HR22, HVKF20, LRF^{+}20, MP13, OF15, WM10]. Sign [ASC^{+}21, BHP19, BCH^{+}21, HAGH21, HZGA08, Hue09, HL10, HGP^{+}17, KLH13, KHE^{+}17, LZL16, LH11, MSDGO19, PSG22, SGOM14, TRLW15, WCH^{+}16]. Sign-language [HAGH21]. Sign-recognition [HAGH21]. Signers [HZGA08]. Signing [AB15]. Signs [AB15]. SIM [TMG23]. Similarities [GGK^{+}20]. Simple [ICL^{+}23]. Simplext [SSB^{+}15]. Simplification [SSB^{+}15]. Simulated [AIB^{+}22, MAB^{+}24]. Site [LRF^{+}20, SHN^{+}11]. Sites [GAY^{+}23, MSDGO19, SGOM14]. Situ [BDKK17]. Situation [FG12]. Situation-Specific [FG12]. Skill [ZAM^{+}21]. Skills [ABLM23, SSG10, TCR^{+}15, WSS^{+}17, WFL24, ZSS^{+}21]. Skin [BDKK17]. Small [SM19]. Social [AED^{+}22, BRH^{+}15, DM23, GAY^{+}23, HKO^{+}23, MS17, SW16, SBPW18, SJPW20, TCR^{+}15, WSS^{+}17]. Socio [SAL^{+}20]. Socio-Technical [SAL^{+}20]. Software [TBC^{+}16, WBO^{+}09]. Solution [LZL16]. Solutions [KKR^{+}23]. Solving [SHN^{+}11]. Someone [FR15]. Sonification [MGBP17, PAD^{+}21, WM10]. Sound [LJYO23, MM24]. Soundscapes [MTM^{+}20]. Source [PK10, RBKC24, VGV^{+}21]. Space [AWBV21, YJGF^{+}22]. Spanish [SSB^{+}15]. Spatial [AWBV21, HR22]. Spatialised [LGG^{+}20]. Spatially [LH11]. Speak [PWH11]. Speakers [MLG^{+}15]. Speaking [RCPA15]. Special [Big17a, Big17b, BOK^{+}22, Bro19, BCH^{+}16, CRPA15, CGH09, FL23, GBK20, Gue23, HS11, HM18a, Hue19, KN22, LL15a, LL15b, MW09, McG12, Mof20, PCRA15, RE24, Ric16a, Ric16b, SAL^{+}20, SH21, Yes12]. Specific [FG12]. Spectrum [BRH^{+}15, BBR^{+}17, MML21, WSS^{+}17, ZFS^{+}18, ZAM^{+}21, ZSS^{+}21]. Speech [APS20, CRPA15, GG16, HBM^{+}21, HVKF20, JW11, KKR^{+}23, LFM15, MLG^{+}15, MST^{+}15, PKL^{+}15, PFM^{+}15, PWH11, PCRA15, RCPA15, RWBM15, VAR^{+}19]. Speech-Language [PWH11]. Speed [Hue09]. Spellcasters [DTD^{+}22]. Spin [TNRW22]. Spindex [JW11]. Spoken [WGMM09]. Spotlights [MTM^{+}20]. Sri [ASZ^{+}21]. SSMI [SMB22]. Staff [WH14]. STANDUP [WBO^{+}09]. Star [ETK18]. State [MPPS23]. Status [FL23, GGK^{+}20, YR19]. steps [LCKG23]. Stop [HAC^{+}15]. Strategies [LHB11, SLHF10, TST^{+}24]. Strategy [WAMG22]. Street [HAC^{+}15]. Stress [MAB^{+}24]. Stroke [APKE11, DTD^{+}22]. Structure [AIP^{+}18]. Student [HGP^{+}17]. Students [MS17]. Studies [PSG22, SVA^{+}18]. Study [APKE11, BRL21, ICL^{+}23, KLH13, LBH23, MEL^{+}22, PHWH14, SAL^{+}20, SW16, VCP^{+}15, WNI^{+}18, BYH12]. Subjects [HKH^{+}09]. Submovements [WG08]. Subspace [MLG^{+}15]. Substitution [KDN23]. Successfully [FR15]. Suffering [PMJB^{+}22]. Suggestions [New08]. Summative [WDEA^{+}21]. Support [BSB22, GRVE20, GBK20, ICL^{+}23, PAD^{+}21]. Supported [FORB^{+}09]. Supporting [ABLM23, HKO^{+}23, JLC^{+}22, KDLD21, LRF^{+}20, RWBM15, WH14]. Surface [BLR^{+}15]. Survey [GF17, LTH^{+}12]. Switch [LCD17]. Symptoms [TST^{+}24]. Syndrome [KDLD21, WFL24, FLKO10, MFKL13]. Synthesis [GAY^{+}23, LH11]. System [GCM15, GGK^{+}20, PHWH14, SSB^{+}15, SMB^{+}19, VCP^{+}15, WAPD08, ZZS^{+}18, ZSSA20, ZAM^{+}21, ZAM^{+}22, ZSS^{+}21]. Systematic [GAY^{+}23, NWVJ22, PSG22]. Systems [BOK^{+}22, PSG22, WGMM09].

Table [BLR^{+}15, WCG^{+}19]. Table-Based [BLR^{+}15]. Tablet [FPC^{+}16, HBM^{+}21]. Tablet-Based [FPC^{+}16, HBM^{+}21]. Tabular [LA22]. TACCESS [BOK^{+}22]. Tactile [BMD^{+}16, Göt18, PWC^{+}22, PBW17, TGMBS24, TMG23, vEPM20]. Tag [MK22]. Tangible [BHH^{+}17, KVFO^{+}23]. Target [HHW13, SM19]. Task [PKGG23, PHWH14, PHB23, SLHF10]. Tasks [CH16, HSKP22, HHMT13, KZM^{+}20, VIH13]. Teach [EGBB23, Sen16, VRL^{+}23]. Teacher [FL23]. Teachers [MS17, PWC^{+}22, PHT^{+}19]. Teaching [CB08, KVFO^{+}23, OBFK15, PDR^{+}16]. TEBRA [PHWH14]. Technical [SAL^{+}20]. Techniques [KVFO^{+}23, MGBP17, OBFK15, PAD^{+}21]. Technologies [AAKS22, BSB22, BBS^{+}21, McG12, PHT^{+}19, RJ10, SSD^{+}22, SW09]. Technology [CRPA15, CH16, CGH09, DAL22, Edw08, FRCH17, GRVE20, GF17, GBK20, KHE^{+}17, MM24, MS17, PCRA15, RCPA15, SW16, TCR^{+}15, Van08, WH14, YMMS22, ORF12]. Technology-Experience [KHE^{+}17]. Teeth [PHWH14]. Telangiectasia [PKGG23]. Telephony [KSV21]. Telepresence [EŠR24]. Telerehabilitation [PMJB^{+}22]. Templates [MSM^{+}15]. Tenets [SBPW18]. Testing [WDEA^{+}21]. Text [EGBB23, GCM15, LYB^{+}22, SSB^{+}15, SC20, SDO^{+}16]. Textual [LJYO23]. Their [AIB^{+}22, AED^{+}22, MM24]. Them [VRL^{+}23]. Thematic [GAY^{+}23]. Therapeutic [MML21]. Therapists [PHT^{+}19]. Therapy [APKE11, HBM^{+}21, LKE^{+}22, PKL^{+}15, PWH11, YMMS22]. there [LCKG23]. Ticket [Sen16]. Time [FM18, MML21]. Tool [LRF^{+}20, AM15, MP13, PKL^{+}15, PRR^{+}22, TNRW22, TFA17]. Tools [AED^{+}22, KVFO^{+}23, LBH23, MPPS23]. Toothbrushing [ZSS^{+}21]. Topicalisation [SFW^{+}24]. Touch [SC20, SM19]. Touch-Enabled [SC20]. Touchscreen [GTU^{+}20, GF17, KZM^{+}20, OBFK15, PG17]. Touchscreen-based [GTU^{+}20]. track [TST^{+}24]. Trackballs [WG08]. Trackers [VGV^{+}21]. Tracking [VIH13]. Tradeoffs [MSDGO19]. Traditional [BGS^{+}16]. Training [AIB^{+}22, SW09, TBB20, ZAM^{+}21]. Trajectories [CB08]. Trajectory [CB08]. Transit [HAC^{+}15]. Transition [LZL16]. Translation [ASC^{+}21]. Translations [BHP19]. Transmitted [TRLW15]. Transparency [MPPS23]. Travel [HR22, KDLD21, vEPM20]. Travelers [GGK^{+}20]. Traveling [MEL^{+}22]. Trends [New08]. Trial [AM15]. Two [DPS^{+}10, SAL^{+}20, SMB22, YJGF^{+}22]. Two-Dimensional [DPS^{+}10]. Two-In-One [YJGF^{+}22]. Type [RBY16]. Typing [NMG^{+}17, PDWT15].

U.S. [AED^{+}22]. Ubiquitous [Van08]. Undergraduates [ZMN^{+}22].

Understandability [KH19]. Understanding [DAL22, LJYO23, LYB^{+}22, MEUHB23, ORF12, PWC^{+}22, PDR^{+}16, PRM23]. Unfamiliar [MEL^{+}22]. Unimanual [YJGF^{+}22]. Universal [WM10]. Universally [SLHF10]. University [PDR^{+}16]. Unreadable [CMT24]. Upper [ETK18, LKRV24]. Urban [HSKP22]. Usability [BHP19, HAGH21, SSG10, SMB^{+}19, WDEA^{+}21]. Usable [SLHF10]. Usage [CH16, FLKO10]. Usages [MEUHB23]. Use [BSB22, DAL22, GGK^{+}20, Her20, JLC^{+}22, LTH^{+}12, MM24, MML21, PHT^{+}19, Sen16, SMB^{+}19, SW09, TST^{+}24, CKA18, ORF12]. Usefulness [VGVG23]. User [GWM^{+}17, ICL^{+}23, JLC^{+}22, BBK^{+}18, LJYO23, MFKL13, MPPS23, NWVJ22, PDWT15, PHWH14, PSG22, PRM23, SVA^{+}18, TMY^{+}09, VCP^{+}15, WAPD08]. Users [ASR^{+}20, CMT24, CB08, DH16, FSR^{+}23, HVKF20, HHMT13, HHW13, KDH21, KVFO^{+}23, LBH23, LHB11, MFKL13, McG12, NMG^{+}17, OF15, PDR^{+}16, SFW^{+}24, SMB^{+}19, SMB22, SJPW20, SGO17, WGMM09, YJGF^{+}22, ZZUK17, vEPM20, SH12]. Using [ATA15, AIP^{+}18, DPS^{+}10, HKH^{+}09, MST^{+}15, MAB^{+}24, MSM^{+}15, PG17, PKGG23, RCW^{+}18, SC20, SHN^{+}11, SDO^{+}16, TCR^{+}15].

validity [BYH12]. VectorEntry [SC20]. Vehicles [BHP^{+}20, FHG21]. Vending [Sen16]. Venues [GRVE20]. Verbs [LH11]. versus [OF15]. Via [CM17, WNI^{+}18, WAMG22]. Vibration [vEPM20]. Video [HKH^{+}09, KSV21, LBH23, LCD17, MSDGO19, SSG10, SGOM14, SW09, TRLW15]. Videos [HGP^{+}17, KLH13, KDH21, SGOM14]. View [HAC^{+}15]. Virtual [AIB^{+}22, AWBV21, ABWV22, BWS^{+}19, BBR^{+}17, DTD^{+}22, ETK18, GWM^{+}17, KDN23, BBK^{+}18, TCR^{+}15, TBB20, TLHC09, VSH^{+}21, ZFS^{+}18, ZZS^{+}18, ZAM^{+}22]. Virtual-world [KDN23]. Visibility [FRCH17]. Vision [APS20, LCKG23, RVL^{+}17, ZSSA20, LA22]. Vision-based [APS20]. Visual [AMB^{+}20, ABLM23, AS10, AWBV21, BHKW22, BLR^{+}15, BHP^{+}20, CMT24, FHG21, GGK^{+}20, GF17, GBK20, KMG^{+}18, LAL^{+}22, LGG^{+}20, MGBP17, MTM^{+}20, MS17, MOL22, MEL^{+}22, PAD^{+}21, SC20, SGO17, SSD^{+}22, TBB20, TLHC09, WNI^{+}18, KDN23]. Visualization [WFL24]. Visualizations [FSR^{+}23, VGVG23]. Visually [AAKS22, CGR^{+}23, CKA18, CLG22, CB08, Göt18, LRF^{+}20, LKE^{+}22, LHB11, MMM^{+}20, PWC^{+}22, SGO17, WM10, ZDW17, ZZUK17]. Vocational [BBR^{+}17]. Voice [ASR^{+}20, ATA15, BMD^{+}16, FORB^{+}09, KSV21, VCP^{+}15]. VR [BOK^{+}22, YJGF^{+}22]. vs [JLC^{+}22, VCP^{+}15].

W4A’22 [RE24]. Walkers [FM18]. Want [MM24]. Wayfinder [AS10]. Ways [FR15]. WCAG [BYH12]. WeAllWalk [FM18]. Wearable [VGV^{+}21]. Weaving [BRDP21]. web [KVH21, FSR^{+}23, KVH21, KVFO^{+}23, LA22, LHB11, MPPS23, MSM^{+}15, RBKC24, STK^{+}10, SHN^{+}11, TMG23, ZZUK17]. Website [BHKW22]. Wheelchair [CCC^{+}16, GMM^{+}15]. Wheelchair-Based [GMM^{+}15]. Whispers [MST^{+}15]. Who [KH19, LKE^{+}22, ZDW17, vEPM20]. Widget [AIP^{+}18]. Wild [SOG^{+}19]. Within [CH16]. Word [AB15, MLG^{+}15, TMY^{+}09]. WordMelodies [ABLM23]. Words [KDH21]. Work [MP13]. Worker [BSB22]. Workers [MSM^{+}15, PHB23]. Working [vEPM20]. Workplace [CLH24]. World [GWM^{+}17, HBM^{+}21, HHMT13, KZM^{+}20, LZL16, RVL^{+}17, TLHC09, KDN23]. Worry [VGV^{+}21]. worse [PRM23]. Write [PWH11]. Write-N-Speak [PWH11].

X [TBB20]. X-Road [TBB20]. XML [DPS^{+}10].

Yoga [RVL^{+}17]. Young [KDLD21, MML21, WFL24]. Younger [MM09].

References

Akter:2022:SPC

[AAKS22] Taslima Akter, Tousif Ahmed, Apu Kapadia, and Manohar Swaminathan. Shared privacy concerns of the visually impaired and sighted bystanders with camera-based assistive technologies. ACM Transactions on Accessible Computing, 15(2):11:1–11:33, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3506857.

Alexanderson:2015:TFA

[AB15] Simon Alexanderson and Jonas Beskow. Towards fully automated motion capture of signs — development and evaluation of a key word signing avatar. ACM Transactions on Accessible Computing, 7(2):7:1–7:??, July 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ahmetovic:2023:WSA

[ABLM23] Dragan Ahmetovic, Cristian Bernareggi, Barbara Leporini, and Sergio Mascetti. WordMelodies: Supporting the acquisition of literacy skills by children with visual impairment through a mobile app. ACM Transactions on Accessible Computing, 16(1):5:1–5:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3565029.

Andrade:2022:PDA

[ABWV22] Ronny Andrade, Steven Baker, Jenny Waycott, and Frank Vetere. A participatory design approach to creating echolocation-enabled virtual environments. ACM Transactions on Accessible Computing, 15(3):18:1–18:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/

doi/10.1145/3516448.

Alonzo:2022:RAT

[AED^{+}22] Oliver Alonzo, Lisa Elliot, Becca Dingman, Sooyeon Lee, Akhter Al Amin, and Matt Huenerfauth. Reading-assistance tools among deaf and hard-of-hearing computing professionals in the U.S.: Their reading experiences, interests and perceptions of social accessibility. ACM Transactions on Accessible Computing, 15 (2):16:1–16:31, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3520198.

Adiani:2022:CIR

[AIB^{+}22] Deeksha Adiani, Aaron Itzkovitz, Dayi Bian, Harrison Katz, Michael Breen, Spencer Hunt, Amy Swanson, Timothy J. Vogus, Joshua Wade, and Nilanjan Sarkar. Career interview readiness in virtual reality (CIRVR): a platform for simulated interview training for autistic individuals and their employers. ACM Transactions on Accessible Computing, 15(1):2:1–2:28, March 2022. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3505560.

Antonelli:2018:DMW

[AIP^{+}18] Humberto Lidio Antonelli, Rodrigo Augusto Igawa, Renata Pontin De Mattos Fortes, Eduardo Henrique Rizo, and Willian Massami Watanabe. Drop-down menu widget identification using HTML structure changes classification. ACM Transactions on Accessible Computing, 11(2):10:1–10:??, June 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Mahmud:2015:IDF

[AM15] Abdullah Al Mahmud and Jean-Bernard Martens. Iterative design and field trial of an aphasia-friendly email tool. ACM Transactions on Accessible Computing, 7(4):13:1–13:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ahmetovic:2020:DLC

[AMB^{+}20] Dragan Ahmetovic, Sergio Mascetti, Cristian Bernareggi, João Guerreiro, Uran Oh, and Chieko Asakawa. Deep learning compensation of rotation errors during navigation assistance for people with visual impairments or blindness. ACM Transactions on Accessible Computing, 12(4):1–19, January 2020. CODEN ????

ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/3349264.

Ahmetovic:2017:MYC

[AMCM17] Dragan Ahmetovic, Roberto Manduchi, James M. Coughlan, and Sergio Mascetti. Mind your crossings: Mining GIS imagery for crosswalk localization. ACM Transactions on Accessible Computing, 9(4):11:1–11:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Allen:2008:FEM

[AMP08] Meghan Allen, Joanna McGrenere, and Barbara Purves. The field evaluation of a mobile digital image communication application designed for people with aphasia. ACM Transactions on Accessible Computing, 1(1):5:1–5:??, May 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Alankus:2011:STT

[APKE11] Gazihan Alankus, Rachel Proffitt, Caitlin Kelleher, and Jack Engsberg. Stroke therapy through motion-based games: a case study. ACM Transactions on Accessible Computing, 4(1):3:1–3:??, November 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ascari:2020:CVB

[APS20] Rúbia E. O. Schultz Ascari, Roberto Pereira, and Luciano Silva. Computer vision-based methodology to improve interaction for people with motor and speech impairment. ACM Transactions on Accessible Computing, 13(4):14:1–14:33, October 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3408300.

Amemiya:2010:OKH

[AS10] Tomohiro Amemiya and Hisashi Sugiyama. Orienting kinesthetically: a haptic handheld wayfinder for people with visual impairments. ACM Transactions on Accessible Computing, 3(2):6:1–6:??, November 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ananthanarayana:2021:DLM

[ASC^{+}21] Tejaswini Ananthanarayana, Priyanshu Srivastava, Akash Chintha, Akhil Santha, Brian Landy, Joseph Panaro, Andre Webster, Nikunj Kotecha, Shagan Sah, Thomastine Sarchet, Raymond Ptucha, and Ifeoma Nwogu. Deep learning methods for sign language translation. ACM Transactions on Accessible Computing, 14(4):22:1–22:30, December 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3477498.

Abdolrahmani:2020:BLS

[ASR^{+}20] Ali Abdolrahmani, Kevin M. Storer, Antony Rishin Mukkath Roy, Ravi Kuber, and Stacy M. Branham. Blind leading the sighted: Drawing design insights from blind users towards more productivity-oriented voice interfaces. ACM Transactions on Accessible Computing, 12(4):1–35, January 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3368426.

Aziz:2022:PYJ

[ASS22] Nida Aziz, Tony Stockman, and Rebecca Stewart. Planning your journey in audio: Design and evaluation of auditory route overviews. ACM Transactions on Accessible Computing, 15(4): 28:1–28:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3531529.

Arachchi:2021:EIS

[ASZ^{+}21] Theja K. Arachchi, Laurianne Sitbon, Jinglan Zhang, Ruwan Gamage, and Priyantha Hewagamage. Enhancing Internet search abilities for people with intellectual disabilities in Sri Lanka. ACM Transactions on Accessible Computing, 14(2):10:1–10:36, July 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3460202.

Aihara:2015:IPV

[ATA15] Ryo Aihara, Tetsuya Takiguchi, and Yasuo Ariki. Individualitypreserving voice conversion for articulation disorders using phoneme-categorized exemplars. ACM Transactions on Accessible Computing, 6(4):13:1–13:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Andrade:2021:EMP

[AWBV21] Ronny Andrade, Jenny Waycott, Steven Baker, and Frank Vetere. Echolocation as a means for people with visual impairment (PVI) to acquire spatial knowledge of virtual space. ACM Transactions on Accessible Computing, 14(1):4:1–4:25, April 2021. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3448273.

Beach:2009:R

[BAS^{+}09] Scott Beach, ACM TACCESS Staff, Richard Schulz, Julie Downs, Judith Matthews, Bruce Barron, and Katherine Seelman. 2008 reviewers. ACM Transactions on Accessible Computing, 2(1):6:1– 6:??, May 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

LalL:2018:EVR

[BBK^{+}18] Lal “Lila” Bozgeyikli, Evren Bozgeyikli, Srinivas Katkoori, Andrew Raij, and Redwan Alqasemi. Effects of virtual reality properties on user experience of individuals with autism. ACM Transactions on Accessible Computing, 11(4):22:1–22:??, November 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3267340.

Bozgeyikli:2017:VRI

[BBR^{+}17] Lal Bozgeyikli, Evren Bozgeyikli, Andrew Raij, Redwan Alqasemi, Srinivas Katkoori, and Rajiv Dubey. Vocational rehabilitation of individuals with autism spectrum disorder with virtual reality. ACM Transactions on Accessible Computing, 10(2):5:1–5:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Bayor:2021:TCB

[BBS^{+}21] Andrew A. Bayor, Margot Brereton, Laurianne Sitbon, Bernd Ploderer, Filip Bircanin, Benoit Favre, and Stewart Koplick. Toward a competency-based approach to co-designing technologies with people with intellectual disability. ACM Transactions on Accessible Computing, 14(2):6:1–6:33, July 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3450355.

Buehler:2016:IIP

[BCH^{+}16] Erin Buehler, Niara Comrie, Megan Hofmann, Samantha McDonald, and Amy Hurst. Investigating the implications of 3D printing in special education. ACM Transactions on Accessible Computing, 8(3):11:1–11:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Bragg:2021:FLS

[BCH^{+}21] Danielle Bragg, Naomi Caselli, Julie A. Hochgesang, Matt Huenerfauth, Leah Katz-Hernandez, Oscar Koller, Raja Kushalnagar, Christian Vogler, and Richard E. Ladner. The FATE landscape of sign language AI datasets: an interdisciplinary perspective. ACM Transactions on Accessible Computing, 14(2):7:1–7:45, July 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/10.1145/3436996.

Betancourt:2017:SCS

[BDKK17] Mariana Aparicio Betancourt, Laura S. Dethorne, Karrie Karahalios, and Jennifer G. Kim. Skin conductance as an in situ marker for emotional arousal in children with neurodevelopmental communication impairments: Methodological considerations and clinical implications. ACM Transactions on Accessible Computing, 9(3):8:1–8:??, February 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Barreto:2009:GE

[BF09] Armando Barreto and Torsten Felzer. Guest editorial. ACM Transactions on Accessible Computing, 2(2):7:1–7:??, June 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Brewer:2016:ETP

[BGS^{+}16] Robin Brewer, Raymundo Cornejo Garcia, Tedmond Schwaba, Darren Gergle, and Anne Marie Piper. Exploring traditional phones as an e-mail interface for older adults. ACM Transactions on Accessible Computing, 8(2):6:1–6:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Baldwin:2017:TDM

[BHH^{+}17] Mark S. Baldwin, Gillian R. Hayes, Oliver L. Haimson, Jennifer Mankoff, and Scott E. Hudson. The tangible desktop: a multimodal approach to nonvisual computing. ACM Transactions on Accessible Computing, 10(3):9:1–9:??, August 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Barbosa:2022:EWP

[BHKW22] Natã M. Barbosa, Jordan Hayes, Smirity Kaushik, and Yang Wang. “Every Website Is a Puzzle!”: Facilitating access to common website features for people with visual impairments. ACM Transactions on Accessible Computing, 15(3):19:1–19:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3519032.

Berke:2019:DPE

[BHP19] Larwan Berke, Matt Huenerfauth, and Kasmira Patel. Design and psychometric evaluation of American Sign Language translations of usability questionnaires. ACM Transactions on Accessible Computing, 12(2):6:1–6:??, July 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3314205.

Brinkley:2020:ENP

[BHP^{+}20] Julian Brinkley, Earl W. Huff, Briana Posadas, Julia Woodward, Shaundra B. Daily, and Juan E. Gilbert. Exploring the needs, preferences, and concerns of persons with visual impairments regarding autonomous vehicles. ACM Transactions on Accessible Computing, 13(1):3:1–3:34, April 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3372280.

Bigham:2017:IASa

[Big17a] Jeffrey Bigham. Introduction to the ASSETS’15 special issue, Part 1. ACM Transactions on Accessible Computing, 9(4):10:1–10:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Bigham:2017:IASb

[Big17b] Jeffrey Bigham. Introduction to the ASSETS’15 special issue, Part 2. ACM Transactions on Accessible Computing, 10(1):1:1– 1:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Binns:2021:HCE

[BK21] Reuben Binns and Reuben Kirkham. How could equality and data protection law shape AI fairness for people with disabilities? ACM Transactions on Accessible Computing, 14(3):17:1–17:32, September 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3473673.

Braier:2015:HSR

[BLR^{+}15] Jonas Braier, Katharina Lattenkamp, Benjamin Räthel, Sandra Schering, Michael Wojatzki, and Benjamin Weyers. Haptic 3D surface representation of table-based data for people with visual impairments. ACM Transactions on Accessible Computing, 6(1): 1:1–1:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Baker:2016:TGV

[BMD^{+}16] Catherine M. Baker, Lauren R. Milne, Ryan Drapeau, Jeffrey Scofield, Cynthia L. Bennett, and Richard E. Ladner. Tactile graphics with a voice. ACM Transactions on Accessible Computing, 8(1):3:1–3:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Biswas:2022:ATS

[BOK^{+}22] Pradipta Biswas, Pilar Orero, Kavita Krishnaswamy, Swaminathan Manohar, and Peter Robinson. ACM TACCESS special issue on adaptive inclusive AR/VR systems. ACM Transactions on Accessible Computing, 15(3):22:1–22:??, September 2022. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3561517.

Borgos-Rodriguez:2021:MDI

[BRDP21] Katya Borgos-Rodriguez, Maitraye Das, and Anne Marie Piper. Melodie: a design inquiry into accessible crafting through audioenhanced weaving. ACM Transactions on Accessible Computing, 14(1):5:1–5:30, April 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3444699.

Boyd:2015:ECI

[BRH^{+}15] Louanne E. Boyd, Kathryn E. Ringland, Oliver L. Haimson, Helen Fernandez, Maria Bistarkey, and Gillian R. Hayes. Evaluating a collaborative iPad game’s impact on social relationships for children with autism spectrum disorder. ACM Transactions on Accessible Computing, 7(1):3:1–3:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Bragg:2021:ELI

[BRL21] Danielle Bragg, Katharina Reinecke, and Richard E. Ladner. Expanding a large inclusive study of human listening rates. ACM Transactions on Accessible Computing, 14(3):12:1–12:26, September 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3461700.

Brock:2019:ISI

[Bro19] Anke Brock. Introduction to the special issue on ASSETS’17. ACM Transactions on Accessible Computing, 12(2):4:1–4:??, July 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/ft_gateway.cfm?id=
3325866.

Balasuriya:2022:SWP

[BSB22] Saminda Sundeepa Balasuriya, Laurianne Sitbon, and Margot Brereton. A support worker perspective on use of new technologies by people with intellectual disabilities. ACM Transactions on Accessible Computing, 15(3):21:1–21:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3523058.

Bian:2019:DPB

[BWS^{+}19] Dayi Bian, Joshua Wade, Amy Swanson, Amy Weitlauf, Zachary Warren, and Nilanjan Sarkar. Design of a physiology-based adaptive virtual reality driving platform for individuals with ASD. ACM Transactions on Accessible Computing, 12(1):2:1– 2:??, February 2019. CODEN ???? ISSN 1936-7228 (print),

1936-7236 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3301498.

Brajnik:2012:ACE

[BYH12] Giorgio Brajnik, Yeliz Yesilada, and Simon Harper. Is accessibility conformance an elusive property? A study of validity and reliability of WCAG 2.0. ACM Transactions on Accessible Computing, 4 (2):8:1–8:??, March 2012. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Crossan:2008:MTP

[CB08] Andrew Crossan and Stephen Brewster. Multimodal trajectory playback for teaching shape information and trajectories to visually impaired computer users. ACM Transactions on Accessible Computing, 1(2):12:1–12:??, October 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Carrington:2016:GRF

[CCC^{+}16] Patrick Carrington, Jian-Ming Chang, Kevin Chang, Catherine Hornback, Amy Hurst, and Shaun K. Kane. The Gest-Rest family: Exploring input possibilities for wheelchair armrests. ACM Transactions on Accessible Computing, 8(3):12:1–12:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Czaja:2009:ISI

[CGH09] Sara J. Czaja, Peter Gregor, and Vicki L. Hanson. Introduction to the special issue on aging and information technology. ACM Transactions on Accessible Computing, 2(1):1:1–1:??, May 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Campos:2023:MGA

[CGR^{+}23] Virgínia P. Campos, Luiz M. G. Gonçalves, Wesnydy L. Ribeiro, Tiago M. U. Araújo, Thaís G. Do Rego, Pedro H. V. Figueiredo, Suanny F. S. Vieira, Thiago F. S. Costa, Caio C. Moraes, Alexandre C. S. Cruz, Felipe A. Araújo, and Guido L. Souza Filho. Machine generation of audio description for blind and visually impaired people. ACM Transactions on Accessible Computing, 16(2):14:1–14:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3590955.

Crabb:2016:AAT

[CH16] Michael Crabb and Vicki L. Hanson. An analysis of age, technology usage, and cognitive characteristics within information retrieval tasks. ACM Transactions on Accessible Computing, 8(3): 10:1–10:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Chakraborty:2018:TDL

[CKA18] Tusher Chakraborty, Taslim Arefin Khan, and A. B. M. Alim Al Islam. Towards devising a low-cost and easy-to-use arithmetic learning framework for economically less-privileged visually impaired children. ACM Transactions on Accessible Computing, 11 (4):21:1–21:??, November 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3265756.

Cooper:2022:DYH

[CLG22] Triston Cooper, Heather Lai, and Jenna Gorlewicz. Do you hear what I hear: The balancing act of designing an electronic hockey puck for playing hockey non-visually. ACM Transactions on Accessible Computing, 15(1):4:1–4:29, March 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3507660.

Coverdale:2024:DAE

[CLH24] Andy Coverdale, Sarah Lewthwaite, and Sarah Horton. Digital accessibility education in context: Expert perspectives on building capacity in academia and the workplace. ACM Transactions on Accessible Computing, 17(2):10:1–10:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3649508.

Cutter:2017:IAM

[CM17] Michael Cutter and Roberto Manduchi. Improving the accessibility of mobile OCR apps via interactive modalities. ACM Transactions on Accessible Computing, 10(4):11:1–11:??, October 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Clarke:2024:IUI

[CMT24] James M. Clarke, Maryam Mehrnezhad, and Ehsan Toreini. Invisible, unreadable, and inaudible cookie notices: an evaluation of cookie notices for users with visual impairments. ACM Transactions on Accessible Computing, 17(1):1:1–1:??, March 2024. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3641281.

Christensen:2015:PSL

[CRPA15] Heidi Christensen, Frank Rudzicz, François Portet, and Jan Alexandersson. Perspectives on speech and language interaction for daily assistive technology: Introduction to Part 1 of the special issue. ACM Transactions on Accessible Computing, 6(3):7:1–7:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Dixon:2022:UHS

[DAL22] Emma Dixon, Jesse Anderson, and Amanda Lazar. Understanding how sensory changes experienced by individuals with a range of age-related cognitive changes can effect technology use. ACM Transactions on Accessible Computing, 15(2):10:1–10:33, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3511906.

Dee:2016:PRU

[DH16] Marianne Dee and Vicki L. Hanson. A pool of representative users for accessibility research: Seeing through the eyes of the users. ACM Transactions on Accessible Computing, 8(1):4:1–4:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Dai:2023:ESS

[DM23] Jiamin Dai and Karyn Moffatt. Enriching social sharing for the dementia community: Insights from in-person and online social programs. ACM Transactions on Accessible Computing, 16 (1):11:1–11:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3582558.

Doush:2010:MPT

[DPS^{+}10] Iyad Abu Doush, Enrico Pontelli, Tran Cao Son, Dominic Simon, and Ou Ma. Multimodal presentation of two-dimensional charts: An investigation using Open Office XML and Microsoft Excel. ACM Transactions on Accessible Computing, 3(2):8:1–8:??, November 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Duval:2022:DSC

[DTD^{+}22] Jared Duval, Rutul Thakkar, Delong Du, Kassandra Chin, Sherry Luo, Aviv Elor, Magy Seif El-Nasr, and Michael John. Designing spellcasters from clinician perspectives: a customizable gesturebased immersive virtual reality game for stroke rehabilitation. ACM Transactions on Accessible Computing, 15(3):26:1–26:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3530820.

Editors:2014:GNE

[Edi14] Editors. Greetings from the new Editors-in-Chief. ACM Transactions on Accessible Computing, 5(3):6:1–6:??, January 2014. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Edwards:2008:KTC

[Edw08] Alistair D. N. Edwards. Keeping up with technology: Commentary on “Computers and People with Disabilities”. ACM Transactions on Accessible Computing, 1(2):8:1–8:??, October 2008. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Edwards:2023:HAT

[EGBB23] Emory J. Edwards, Michael Gilbert, Emily Blank, and Stacy M. Branham. How the Alt text gets made: What roles and processes of Alt text creation can teach us about inclusive imagery.

ACM Transactions on Accessible Computing, 16(2):18:1–18:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3587469.

Elmimouni:2024:NCC

[EŠR24] Houda Elmimouni, Selma Šabanović, and Jennifer A. Rode. Navigating the cyborg classroom: Telepresence robots, accessibility challenges, and inclusivity in the classroom. ACM Transactions on Accessible Computing, 17(2):8:1–8:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3672569.

Elor:2018:PSC

[ETK18] Aviv Elor, Mircea Teodorescu, and Sri Kurniawan. Project Star Catcher: a novel immersive virtual reality experience for upper limb rehabilitation. ACM Transactions on Accessible Computing, 11(4):20:1–20:??, November 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3265755.

Flatla:2012:SSM

[FG12] David R. Flatla and Carl Gutwin. Situation-specific models of color differentiation. ACM Transactions on Accessible Computing, 4(3):13:1–13:??, December 2012. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Fink:2021:FAV

[FHG21] Paul D. S. Fink, Jessica A. Holz, and Nicholas A. Giudice. Fully autonomous vehicles for people with visual impairment: Policy, accessibility, and future directions. ACM Transactions on Accessible Computing, 14(3):15:1–15:17, September 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3471934.

Forster:2023:DMI

[FL23] Andreas Förster and Steffen Lepa. Digital musical instruments in special educational needs schools: Requirements from the music teachers’ perspective and the status quo in Germany. ACM Transactions on Accessible Computing, 16(3):23:1–23:??, September 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3616015.

Feng:2010:CUC

[FLKO10] Jinjuan Feng, Jonathan Lazar, Libby Kumin, and Ant Ozok. Computer usage by children with Down Syndrome: Challenges and future research. ACM Transactions on Accessible Computing, 2(3): 13:1–13:??, March 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Flores:2018:WAD

[FM18] Germán H. Flores and Roberto Manduchi. WeAllWalk: an annotated dataset of inertial sensor time series from blind walkers. ACM Transactions on Accessible Computing, 11(1):4:1–4:??, April 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Fried-Oken:2009:EVO

[FORB^{+}09] Melanie Fried-Oken, Charity Rowland, Glory Baker, Mayling Dixon, Carolyn Mills, Darlene Schultz, and Barry Oken. The effect of voice output on AAC-supported conversations of persons with Alzheimer’s disease. ACM Transactions on Accessible Computing, 1(3):15:1–15:??, February 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Fage:2016:TBA

[FPC^{+}16] Charles Fage, Léonard Pommereau, Charles Consel, Emilie Balland, and Hélène Sauzéon. Tablet-based activity schedule in mainstream environment for children with autism and children with ID. ACM Transactions on Accessible Computing, 8(3):9:1–9:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Felzer:2015:ESN

[FR15] Torsten Felzer and Stephan Rinderknecht. Experiences of someone with a neuromuscular disease in operating a PC (and ways to successfully overcome challenges). ACM Transactions on Accessible Computing, 6(2):4:1–4:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Faucett:2017:VDA

[FRCH17] Heather A. Faucett, Kate E. Ringland, Amanda L. L. Cullen, and Gillian R. Hayes. (In)Visibility in disability and assistive technology. ACM Transactions on Accessible Computing, 10(4):14:1– 14:??, October 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Fan:2023:ADV

[FSR^{+}23] Danyang Fan, Alexa Fay Siu, Hrishikesh Rao, Gene Sung-Ho Kim, Xavier Vazquez, Lucy Greco, Sile O’Modhrain, and Sean Follmer. The accessibility of data visualizations on the Web for screen reader users: Practices and experiences during COVID-19. ACM Transactions on Accessible Computing, 16(1):4:1–4:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3557899.

Fan:2018:GCS

[FT18] Mingming Fan and Khai N. Truong. Guidelines for creating seniorfriendly product instructions. ACM Transactions on Accessible Computing, 11(2):9:1–9:??, June 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ghenai:2023:GDS

[GAY^{+}23] Amira Ghenai, Philips Ayeni, Jing Yu, Robin Cohen, and Karyn Moffatt. Guidelines for designing social networking sites for older adults: a systematic review with thematic synthesis. ACM Transactions on Accessible Computing, 16(3):19:1–19:??, September 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3615662.

Guerreiro:2020:ISI

[GBK20] João Guerreiro, Anke M. Brock, and Hernisa Kacorri. Introduction to the special issue on technology to support independent orientation and mobility of people with visual impairments. ACM Transactions on Accessible Computing, 13(2):5e:1– 5e:2, June 2020. CODEN ???? ISSN 1936-7228 (print), 1936-

7236 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3398652.

Garcia:2015:MPL

[GCM15] Luís Filipe Garcia, Luís Caldas De Oliveira, and David Martins De Matos. Measuring the performance of a location-aware text prediction system. ACM Transactions on Accessible Computing, 7(1):2:1–2:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Grussenmeyer:2017:ATT

[GF17] William Grussenmeyer and Eelke Folmer. Accessible touchscreen technology for people with visual impairments: a survey. ACM Transactions on Accessible Computing, 9(2):6:1–6:??, January 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Guerreiro:2016:SDC

[GG16] João Guerreiro and Daniel Gonçalves. Scanning for digital content: How blind and sighted people perceive concurrent speech. ACM Transactions on Accessible Computing, 8(1):2:1–2:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Giudice:2020:UIN

[GGK^{+}20] Nicholas A. Giudice, Benjamin A. Guenther, Toni M. Kaplan, Shane M. Anderson, Robert J. Knuesel, and Joseph F. Cioffi. Use of an indoor navigation system by sighted and blind travelers: Performance similarities across visual status and age. ACM Transactions on Accessible Computing, 13(3):11:1–11:27, August 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/10.1145/3407191.

Guerreiro:2020:GNE

[GL20] Tiago Guerreiro and Stephanie Ludi. Greetings from the new Editors-in-Chief. ACM Transactions on Accessible Computing, 12 (4):1, January 2020. CODEN ???? ISSN 1936-7228 (print), 1936-

7236 (electronic). URL https://dl.acm.org/doi/abs/10.1145/
3372922.

Guerreiro:2021:EEC

[GL21] Tiago Guerreiro and Stephanie Ludi. Editorial from the Editorsin-Chief. ACM Transactions on Accessible Computing, 14(1):1:1– 1:2, April 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3456772.

Gerling:2015:DWB

[GMM^{+}15] Kathrin M. Gerling, Regan L. Mandryk, Matthew Miller, Michael R. Kalyn, Max Birk, and Jan D. Smeddinck. Designing wheelchair-based movement games. ACM Transactions on Accessible Computing, 6(2):6:1–6:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Gotzelmann:2018:VAA

[Göt18] T. Götzelmann. Visually augmented audio-tactile graphics for visually impaired people. ACM Transactions on Accessible Computing, 11(2):8:1–8:??, June 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ghedira:2009:CEC

[GPB09] Souhir Ghedira, Pierre Pino, and Guy Bourhis. Conception and experimentation of a communication device with adaptive scanning. ACM Transactions on Accessible Computing, 1(3):14:1– 14:??, February 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Gerling:2020:CRT

[GRVE20] Kathrin Gerling, Mo Ray, Vero Vanden Abeele, and Adam B. Evans. Critical reflections on technology to support physical activity among older adults: an exploration of leading HCI venues. ACM Transactions on Accessible Computing, 13(1):1:1–1:23, April 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3374660.

Gorlewicz:2020:DGR

[GTU^{+}20] Jenna L. Gorlewicz, Jennifer L. Tennison, P. Merlin Uesbeck, Margaret E. Richard, Hari P. Palani, Andreas Stefik, Derrick W. Smith, and Nicholas A. Giudice. Design guidelines and recommendations for multimodal, touchscreen-based graphics. ACM Transactions on Accessible Computing, 13(3):10:1–10:30, August 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/10.1145/3403933.

Guerreiro:2023:ISI

[Gue23] João Guerreiro. Introduction to the special issue on ASSETS’21. ACM Transactions on Accessible Computing, 16(2):15:1–15:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3605947.

Galliers:2017:EEP

[GWM^{+}17] Julia Galliers, Stephanie Wilson, Jane Marshall, Richard Talbot, Niamh Devane, Tracey Booth, Celia Woolf, and Helen Greenwood. Experiencing EVA Park, a multi-user virtual world for people with aphasia. ACM Transactions on Accessible Computing, 10(4):15:1– 15:??, October 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Glinert:2008:CPD

[GY08] Ephraim P. Glinert and Bryant W. York. Computers and people with disabilities. ACM Transactions on Accessible Computing, 1 (2):7:1–7:??, October 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hara:2015:IPT

[HAC^{+}15] Kotaro Hara, Shiri Azenkot, Megan Campbell, Cynthia L. Bennett, Vicki Le, Sean Pannella, Robert Moore, Kelly Minckler, Rochelle H. Ng, and Jon E. Froehlich. Improving public transit accessibility for blind riders by crowdsourcing bus stop landmark locations with Google Street View: an extended analysis. ACM Transactions on Accessible Computing, 6(2):5:1–5:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hassan:2021:ESR

[HAGH21] Saad Hassan, Oliver Alonzo, Abraham Glasser, and Matt Huenerfauth. Effect of sign-recognition performance on the usability of sign-language dictionary search. ACM Transactions on Accessible Computing, 14(4):18:1–18:33, December 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3470650.

Hair:2021:LET

[HBM^{+}21] Adam Hair, Kirrie J. Ballard, Constantina Markoulli, Penelope Monroe, Jacqueline Mckechnie, Beena Ahmed, and Ricardo Gutierrez-Osuna. A longitudinal evaluation of tablet-based child speech therapy with apraxia world. ACM Transactions on Accessible Computing, 14(1):3:1–3:26, April 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3433607.

Hersh:2020:MMU

[Her20] Marion Hersh. Mental maps and the use of sensory information by blind and partially sighted people. ACM Transactions on Accessible Computing, 13(2):6:1–6:32, June 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3375279.

Hu:2015:IIS

[HF15] Ruimin Hu and Jinjuan Heidi Feng. Investigating information search by people with cognitive disabilities. ACM Transactions on Accessible Computing, 7(1):1:1–1:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Huenerfauth:2017:ELF

[HGP^{+}17] Matt Huenerfauth, Elaine Gale, Brian Penly, Sree Pillutla, Mackenzie Willard, and Dhananjai Hariharan. Evaluation of language feedback methods for student videos of American Sign Language. ACM Transactions on Accessible Computing, 10(1):2:1– 2:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hurst:2013:DUP

[HHMT13] Amy Hurst, Scott E. Hudson, Jennifer Mankoff, and Shari Trewin. Distinguishing users by pointing performance in laboratory and real-world tasks. ACM Transactions on Accessible Computing, 5 (2):5:1–5:??, October 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hwang:2013:ETE

[HHW13] Faustina Hwang, Nic Hollinworth, and Nitin Williams. Effects of target expansion on selection performance in older computer users. ACM Transactions on Accessible Computing, 5(1):1:1–1:??, September 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hailpern:2009:AHC

[HKH^{+}09] Joshua Hailpern, Karrie Karahalios, James Halle, Laura Dethorne, and Mary-Kelsey Coletto. A3: HCI coding guideline for research using video annotation to assess behavior of nonverbal subjects with computer-based intervention. ACM Transactions on Accessible Computing, 2(2):8:1–8:??, June 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hamidi:2023:SSI

[HKO^{+}23] Foad Hamidi, Tsion Kidane, Patrick Mbullo Owuor, Michaela Hynie, and Melanie Baljko. Supporting social inclusion with DIY-ATs: Perspectives of Kenyan caregivers of children with cognitive disabilities. ACM Transactions on Accessible Computing, 16 (3):20:1–20:??, September 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3616378.

Huenerfauth:2010:AAM

[HL10] Matt Huenerfauth and Pengfei Lu. Accurate and accessible motion-capture glove calibration for sign language data collection. ACM Transactions on Accessible Computing, 3(1):2:1–2:??, September 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Huenerfauth:2018:IAS

[HM18a] Matt Huenerfauth and Kathleen F. McCoy. Introduction to the ASSETS’16 special issue. ACM Transactions on Accessible Computing, 11(1):1:1–1:??, April 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Huenerfauth:2018:II

[HM18b] Matt Huenerfauth and Kathleen F. McCoy. Introduction to this issue. ACM Transactions on Accessible Computing, 11(2):7:1–7:??, June 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Huenerfauth:2019:EMO

[HM19] Matt Huenerfauth and Kathleen F. McCoy. Editorial: a message from the outgoing Editors-in-Chief. ACM Transactions on Accessible Computing, 12(3):10:1–10:??, September 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3345019.

Hersh:2022:RDS

[HR22] Marion Hersh and Alejandro Rafael Garcia Ramirez. Route descriptions, spatial knowledge and spatial representations of blind and partially sighted people: Improved design of electronic travel aids. ACM Transactions on Accessible Computing, 15(4): 32:1–32:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3549077.

Hanson:2011:IAS

[HS11] Vicki L. Hanson and Andrew Sears. Introduction ASSETS’10 special issue. ACM Transactions on Accessible Computing, 4(1):1:1– 1:??, November 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hanson:2012:IA

[HS12] Vicki L. Hanson and Andrew Sears. Introduction to article 7. ACM Transactions on Accessible Computing, 4(2):6:1–6:??, March 2012. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Hoogsteen:2022:BCD

[HSKP22] Karst M. P. Hoogsteen, Sarit Szpiro, Gabriel Kreiman, and Eli Peli. Beyond the cane: Describing urban scenes to blind people for mobility tasks. ACM Transactions on Accessible Computing, 15(3):20:1–20:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3522757.

Huenerfauth:2009:LMM

[Hue09] Matt Huenerfauth. A linguistically motivated model for speed and pausing in animations of American Sign Language. ACM Transactions on Accessible Computing, 2(2):9:1–9:??, June 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Huenerfauth:2019:ISI

[Hue19] Kathleen F. McCoy / Matt Huenerfauth. Introduction to the special issue on ASSETS’17 (part 2). ACM Transactions on Accessible Computing, 12(3):13:1–13:??, September 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3345021.

Hong:2020:RSI

[HVKF20] Jonggi Hong, Christine Vaing, Hernisa Kacorri, and Leah Findlater. Reviewing speech input with audio: Differences between blind and sighted users. ACM Transactions on Accessible Computing, 13(1):2:1–2:28, April 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3382039.

Huenerfauth:2008:EAS

[HZGA08] Matt Huenerfauth, Liming Zhao, Erdan Gu, and Jan Allbeck. Evaluation of American Sign Language generation by native ASL signers. ACM Transactions on Accessible Computing, 1(1):3:1– 3:??, May 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Iniesto:2023:CSC

[ICL^{+}23] Francisco Iniesto, Tim Coughlan, Kate Lister, Peter Devine, Nick Freear, Richard Greenwood, Wayne Holmes, Ian Kenny, Kevin McLeod, and Ruth Tudor. Creating ‘a simple conversation’: Designing a conversational user interface to improve the experience of accessing support for study. ACM Transactions on Accessible Computing, 16(1):6:1–6:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3568166.

Ilsar:2022:IIE

[IKTF22] Alon Ilsar, Gail Kenning, Sam Trolland, and Ciaran Frame. Inclusive improvisation: Exploring the line between listening and playing music. ACM Transactions on Accessible Computing, 15(2):9:1– 9:21, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3506856.

Jamieson:2022:SPA

[JLC^{+}22] Matthew Jamieson, Marilyn Lennon, Breda Cullen, Stephen Brewster, and Jonathan Evans. Supporting people with acquired brain injury to use a reminding app; narrow-deep vs. broadshallow user interfaces. ACM Transactions on Accessible Computing, 15(1):1:1–1:23, March 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3501275.

Jeon:2011:SSI

[JW11] Myounghoon Jeon and Bruce N. Walker. Spindex (speech index) improves auditory menu acceptance and navigation performance. ACM Transactions on Accessible Computing, 3(3):10:1– 10:??, April 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Kafle:2021:DHH

[KDH21] Sushant Kafle, Becca Dingman, and Matt Huenerfauth. Deaf and hard-of-hearing users evaluating designs for highlighting key words in educational lecture videos. ACM Transactions on Accessible Computing, 14(4):20:1–20:24, December 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3470651.

Khan:2021:TDM

[KDLD21] Al Majed Khan, Dr Mark D. Dunlop, Dr Marilyn Lennon, and Dr Mateusz Dubiel. Towards designing mobile apps for independent travel: Exploring current barriers and opportunities for supporting young adults with Down’s syndrome. ACM Transactions on Accessible Computing, 14(3):13:1–13:40, September 2021. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3460943.

Kruger:2023:MCB

[KDN23] Rynhardt Kruger, Febe De Wet, and Thomas Niesler. Mathematical content browsing for print-disabled readers based on virtual-world exploration and audio-visual sensory substitution. ACM Transactions on Accessible Computing, 16(2):12:1–12:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3584365.

Kafle:2019:PUI

[KH19] Sushant Kafle and Matt Huenerfauth. Predicting the understandability of imperfect English captions for people who are deaf or hard of hearing. ACM Transactions on Accessible Computing, 12 (2):7:1–7:??, July 2019. CODEN ???? ISSN 1936-7228 (print),

1936-7236 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3325862.

Kacorri:2017:RAD

[KHE^{+}17] Hernisa Kacorri, Matt Huenerfauth, Sarah Ebling, Kasmira Patel, Kellie Menzies, and Mackenzie Willard. Regression analysis of demographic and technology-experience factors influencing acceptance of sign language animation. ACM Transactions on Accessible Computing, 10(1):3:1–3:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Kuhn:2023:MAA

[KKR^{+}23] Korbinian Kuhn, Verena Kersken, Benedikt Reuter, Niklas Egger, and Gottfried Zimmermann. Measuring the accuracy of automatic speech recognition solutions. ACM Transactions on Accessible Computing, 16(4):25:1–25:??, December 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3636513.

Kushalnagar:2014:AEC

[KLB14] Raja S. Kushalnagar, Walter S. Lasecki, and Jeffrey P. Bigham. Accessibility evaluation of classroom captions. ACM Transactions on Accessible Computing, 5(3):7:1–7:??, January 2014. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Kacorri:2013:EDH

[KLH13] Hernisa Kacorri, Pengfei Lu, and Matt Huenerfauth. Effect of displaying human videos during an evaluation study of American Sign Language animation. ACM Transactions on Accessible Computing, 5(2):4:1–4:??, October 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Kacorri:2018:IAO

[KMG^{+}18] Hernisa Kacorri, Sergio Mascetti, Andrea Gerino, Dragan Ahmetovic, Valeria Alampi, Hironobu Takagi, and Chieko Asakawa. Insights on assistive orientation and mobility of people with visual impairment based on large-scale longitudinal data. ACM Transactions on Accessible Computing, 11(1):5:1–5:??, April 2018. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Kouroupetroglou:2022:ISI

[KN22] Georgios Kouroupetroglou and Hugo Nicolau. Introduction to the special issue on ASSETS’20. ACM Transactions on Accessible Computing, 15(2):8e:1–8e:2, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/

doi/10.1145/3542810.

Kozma-Spytek:2021:FAA

[KSV21] Linda Kozma-Spytek and Christian Vogler. Factors affecting the accessibility of voice telephony for people with hearing loss: Audio encoding, network impairments, video and environmental noise. ACM Transactions on Accessible Computing, 14(4):21:1–21:35, December 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3479160.

Kearney-Volpe:2023:TPT

[KVFO^{+}23] Claire Kearney-Volpe, Chancey Fleet, Keita Ohshiro, Veronica Alfaro Arias, Eric Hao Xu, and Amy Hurst. Tangible progress: Tools, techniques, and impacts of teaching Web development to screen reader users. ACM Transactions on Accessible Computing, 16(1):8:1–8:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3585315.

Kearney-Volpe:2021:AWD

[KVH21] Claire Kearney-Volpe and Amy Hurst. Accessible Web development: Opportunities to improve the education and practice of web development with a screen reader. ACM Transactions on Accessible Computing, 14(2):8:1–8:32, July 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3458024.

Kuber:2011:EHH

[KYO11] Ravi Kuber, Wai Yu, and M. Sile O’Modhrain. Evaluation of haptic HTML mappings derived from a novel methodology. ACM Transactions on Accessible Computing, 3(4):12:1–12:??, April 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Kim:2020:CTM

[KZM^{+}20] Sung-Hee Kim, Kailun Zhang, Joanna McGrenere, Kellogg S. Booth, and Claudia Jacova. A comparison of touchscreen and mouse for real-world and abstract tasks with older adults. ACM Transactions on Accessible Computing, 13(4):16:1–16:26, October 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3418057.

Lee:2022:CTA

[LA22] Hae-Na Lee and Vikas Ashok. Customizable tabular access to Web data records for convenient low-vision screen magnifier interaction. ACM Transactions on Accessible Computing, 15(2):14:1–14:22, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3517044.

Ladner:2008:AEC

[Lad08] Richard E. Ladner. Access and empowerment: Commentary on “Computers and People with Disabilities”. ACM Transactions on Accessible Computing, 1(2):11:1–11:??, October 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Lee:2022:AAR

[LAL^{+}22] Sooyeon Lee, Nelson Daniel Troncoso Aldas, Chonghan Lee, Mary Beth Rosson, John M. Carroll, and Vijaykrishnan Narayanan. AIGuide: Augmented reality hand guidance in a visual prosthetic. ACM Transactions on Accessible Computing, 15(2):12:1–12:32, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3508501.

Leporini:2023:VCT

[LBH23] Barbara Leporini, Marina Buzzi, and Marion Hersh. Video conferencing tools: Comparative study of the experiences of screen reader users and the development of more inclusive design guidelines. ACM Transactions on Accessible Computing, 16(1):7:1–7:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3573012.

Lopez:2017:DDO

[LCD17] Sebastián Aced López, Fulvio Corno, and Luigi De Russis. Design and development of one-switch video games for children with severe motor disabilities. ACM Transactions on Accessible Computing, 10(4):12:1–12:??, October 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Lu:2023:WTM

[LCKG23] Leon Lu, Karen Anne Cochrane, Jin Kang, and Audrey Girouard. “Why are there so many steps?”: Improving access to blind and low vision music learning through personal adaptations and future design ideas. ACM Transactions on Accessible Computing, 16(3):22:1–22:??, September 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3615663.

Leung:2010:MLI

[LFM^{+}10] Rock Leung, Leah Findlater, Joanna McGrenere, Peter Graf, and Justine Yang. Multi-layered interfaces to improve older adults’ initial learnability of mobile applications. ACM Transactions on Accessible Computing, 3(1):1:1–1:??, September 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Laaridh:2015:ADP

[LFM15] Imed Laaridh, Corinne Fredouille, and Christine Meunier. Automatic detection of phone-based anomalies in dysarthric speech. ACM Transactions on Accessible Computing, 6(3):9:1–9:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Lock:2020:EAS

[LGG^{+}20] Jacobus C. Lock, Iain D. Gilchrist, Iain D. Gilchrist, Grzegorz Cielniak, and Nicola Bellotto. Experimental analysis of a spatialised audio interface for people with visual impairments. ACM Transactions on Accessible Computing, 13(4):17:1–17:21, October 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3412325.

Lu:2011:DDS

[LH11] Pengfei Lu and Matt Huenerfauth. Data-driven synthesis of spatially inflected verbs for American Sign Language animation. ACM Transactions on Accessible Computing, 4(1):4:1–4:??, November 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Lunn:2011:IBS

[LHB11] Darren Lunn, Simon Harper, and Sean Bechhofer. Identifying behavioral strategies of visually impaired users to improve access to Web content. ACM Transactions on Accessible Computing, 3 (4):13:1–13:??, April 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Lee:2023:AUU

[LJYO23] Yun Jung Lee, Hwayeon Joh, Suhyeon Yoo, and Uran Oh. AccessComics2: Understanding the user experience of an accessible comic book reader for blind people with textual sound effects. ACM Transactions on Accessible Computing, 16(1):2:1–2:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3555720.

Li:2022:SRA

[LKE^{+}22] Yifan Li, Kangsoo Kim, Austin Erickson, Nahal Norouzi, Jonathan Jules, Gerd Bruder, and Gregory F. Welch. A scoping review of assistance and therapy with head-mounted displays for people who are visually impaired. ACM Transactions on Accessible Computing, 15(3):25:1–25:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3522693.

Lewis:2024:AIA

[LKRV24] Brittany Lewis, Priyankan Kirupaharan, Tina-Marie Ranalli, and Krishna Venkatasubramanian. A3C: an image-association-based computing device authentication framework for people with upper extremity impairments. ACM Transactions on Accessible Computing, 17(2):6:1–6:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3652522.

Ladner:2015:IASa

[LL15a] Richard Ladner and Jonathan Lazar. Introduction to the AS-SETS’13 special issue. ACM Transactions on Accessible Computing, 6(2):4:1–4:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ladner:2015:IASb

[LL15b] Richard Ladner and Jonathan Lazar. Introduction to the AS-SETS’13 special issue, Part 2. ACM Transactions on Accessible Computing, 7(3):8:1–8:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Leporini:2020:DGI

[LRF^{+}20] Barbara Leporini, Valentina Rossetti, Francesco Furfari, Susanna Pelagatti, and Andrea Quarta. Design guidelines for an interactive 3D model as a supporting tool for exploring a cultural site by visually impaired and sighted people. ACM Transactions on Accessible Computing, 13(3):9:1–9:39, August 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3399679.

Leung:2012:HOA

[LTH^{+}12] Rock Leung, Charlotte Tang, Shathel Haddad, Joanna Mcgrenere, Peter Graf, and Vilia Ingriany. How older adults learn to use mobile devices: Survey and field investigations. ACM Transactions on Accessible Computing, 4(3):11:1–11:??, December 2012. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Lottridge:2022:AUT

[LYB^{+}22] Danielle Lottridge, Chris Yoon, Darren Burton, Chester Wang, and Jofish Kaye. Ally: Understanding text messaging to build a better onscreen keyboard for blind people. ACM Transactions on Accessible Computing, 15(4):30:1–30:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3533707.

Li:2016:STM

[LZL16] Kehuang Li, Zhengyu Zhou, and Chin-Hui Lee. Sign transition modeling and a scalable solution to continuous sign language recognition for real-world applications. ACM Transactions on Accessible Computing, 8(2):7:1–7:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Migovich:2024:SDA

[MAB^{+}24] Miroslava Migovich, Deeksha Adiani, Michael Breen, Amy Swanson, Timothy J. Vogus, and Nilanjan Sarkar. Stress detection of autistic adults during simulated job interviews using a novel physiological dataset and machine learning. ACM Transactions on Accessible Computing, 17(1):2:1–2:??, March 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3639709.

McCoy:2010:GEA

[McC10] Kathleen F. McCoy. Guest editorial: ASSETS 2009. ACM Transactions on Accessible Computing, 3(2):4:1–4:??, November 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

McGrenere:2012:ISI

[McG12] Joanna McGrenere. Introduction to special issue on mobile technologies for older users. ACM Transactions on Accessible Computing, 4(3):10:1–10:??, December 2012. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Muller:2022:TMI

[MEL^{+}22] Karin Müller, Christin Engel, Claudia Loitsch, Rainer Stiefelhagen, and Gerhard Weber. Traveling more independently: a study on the diverse needs and challenges of people with visual or mobility impairments in unfamiliar indoor environments. ACM Transactions on Accessible Computing, 15(2):13:1–13:44, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/10.1145/3514255.

Momotaz:2023:UUL

[MEUHB23] Farhani Momotaz, Md Ehtesham-Ul-Haque, and Syed Masum Billah. Understanding the usages, lifecycle, and opportunities of screen readers’ plugins. ACM Transactions on Accessible Computing, 16(2):17:1–17:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3582697.

Ma:2013:IUB

[MFKL13] Yao Ma, Jinjuan Feng, Libby Kumin, and Jonathan Lazar. Investigating user behavior for authentication methods: a comparison between individuals with Down Syndrome and neurotypical users. ACM Transactions on Accessible Computing, 4(4):15:1–15:??, July 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Mascetti:2017:ENS

[MGBP17] Sergio Mascetti, Andrea Gerino, Cristian Bernareggi, and Lorenzo Picinali. On the evaluation of novel sonification techniques for non-visual shape exploration. ACM Transactions on Accessible Computing, 9(4):13:1–13:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Muehlbradt:2022:WAT

[MK22] Annika Muehlbradt and Shaun K. Kane. What’s in an ALT tag? Exploring caption content priorities through collaborative captioning. ACM Transactions on Accessible Computing, 15(1):6:1–6:32, March 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3507659.

Martinez:2015:IAS

[MLG^{+}15] David Martínez, Eduardo Lleida, Phil Green, Heidi Christensen, Alfonso Ortega, and Antonio Miguel. Intelligibility assessment and speech recognizer word accuracy rate prediction for dysarthric speakers in a factor analysis subspace. ACM Transactions on Accessible Computing, 6(3):10:1–10:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Moffatt:2009:EMI

[MM09] Karyn Moffatt and Joanna McGrenere. Exploring methods to improve pen-based menu selection for younger and older adults. ACM Transactions on Accessible Computing, 2(1):3:1–3:??, May 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Marsh:2024:DWS

[MM24] Abigail Marsh and Lauren R. Milne. I don’t want to sound rude, but it’s none of their business: Exploring security and privacy concerns around assistive technology use in educational settings. ACM Transactions on Accessible Computing, 17(2):7:1–7:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3670690.

Mcgowan:2021:EUR

[MML21] John J. Mcgowan, Iain Mcgregor, and Gregory Leplatre. Evaluation of the use of real-time 3D graphics to augment therapeutic music sessions for young people on the autism spectrum. ACM Transactions on Accessible Computing, 14(1):2:1–2:41, April 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/10.1145/3445032.

Madugalla:2020:CAO

[MMM^{+}20] Anuradha Madugalla, Kim Marriott, Simone Marinai, Samuele Capobianco, and Cagatay Goncu. Creating accessible online floor plans for visually impaired readers. ACM Transactions on Accessible Computing, 13(4):15:1–15:37, October 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3410446.

Moffatt:2020:ISI

[Mof20] Karyn Moffatt. Introduction to the special issue on ASSETS’18. ACM Transactions on Accessible Computing, 12(4):1, January 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/abs/10.1145/3372925.

Mountapmbeme:2022:AAB

[MOL22] Aboubakar Mountapmbeme, Obianuju Okafor, and Stephanie Ludi. Addressing accessibility barriers in programming for people with visual impairments: a literature review. ACM Transactions on Accessible Computing, 15(1):7:1–7:26, March 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507469.

Moll:2013:HTG

[MP13] Jonas Moll and Eva-Lotta Sallnäs Pysander. A haptic tool for group work on geometrical concepts engaging blind and sighted pupils. ACM Transactions on Accessible Computing, 4(4):14:1– 14:??, July 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Manca:2023:TAW

[MPPS23] Marco Manca, Vanessa Palumbo, Fabio Paternò, and Carmen Santoro. The transparency of automatic Web accessibility evaluation tools: Design criteria, state of the art, and user perception. ACM Transactions on Accessible Computing, 16(1):3:1–3:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3556979.

Morash:2017:SPA

[MS17] Valerie S. Morash and Yue-Ting Siu. Social predictors of assistive technology proficiency among teachers of students with visual impairments. ACM Transactions on Accessible Computing, 9(2): 4:1–4:??, January 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Monteiro:2019:TED

[MSDGO19] Caio D. D. Monteiro, Frank M. Shipman, Satyakiran Duggina, and Ricardo Gutierrez-Osuna. Tradeoffs in the efficient detection of sign language content in video sharing sites. ACM Transactions on Accessible Computing, 12(2):5:1–5:??, July 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/ft_gateway.cfm?id=3325863.

Morash:2015:GNW

[MSM^{+}15] Valerie S. Morash, Yue-Ting Siu, Joshua A. Miele, Lucia Hasty, and Steven Landau. Guiding novice Web workers in making image descriptions using templates. ACM Transactions on Accessible Computing, 7(4):12:1–12:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Mcloughlin:2015:RPS

[MST^{+}15] Ian V. Mcloughlin, Hamid Reza Sharifzadeh, Su Lim Tan, Jingjie Li, and Yan Song. Reconstruction of phonated speech from whispers using formant-derived plausible pitch modulation. ACM Transactions on Accessible Computing, 6(4):12:1–12:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Menzies:2022:ARC

[MTC22] Rachel Menzies, Garreth W. Tigwell, and Michael Crabb. Author reflections on creating accessible academic papers. ACM Transactions on Accessible Computing, 15(4):33:1–33:??, December 2022.

CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/10.1145/3546195.

May:2020:SSD

[MTM^{+}20] Keenan R. May, Brianna J. Tomlinson, Xiaomeng Ma, Phillip Roberts, and Bruce N. Walker. Spotlights and soundscapes: On the design of mixed reality auditory environments for persons with visual impairment. ACM Transactions on Accessible Computing, 13(2):8:1–8:47, June 2020. CODEN ???? ISSN 1936-7228 (print),

1936-7236 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3378576.

McCoy:2009:ISI

[MW09] Kathleen F. McCoy and Annalu Waller. Introduction to the special issue on AAC. ACM Transactions on Accessible Computing, 1(3):13:1–13:??, February 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Newell:2008:ACP

[New08] Alan F. Newell. Accessible computing — past trends and future suggestions: Commentary on “Computers and People with Disabilities”. ACM Transactions on Accessible Computing, 1(2): 9:1–9:??, October 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Nicolau:2017:ILE

[NMG^{+}17] Hugo Nicolau, Kyle Montague, Tiago Guerreiro, André Rodrigues, and Vicki L. Hanson. Investigating laboratory and everyday typing performance of blind users. ACM Transactions on Accessible Computing, 10(1):4:1–4:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Nolte:2022:IAB

[NWVJ22] Amelie Nolte, Jacob Wobbrock, Torben Volkmann, and Nicole Jochems. Implementing ability-based design: a systematic approach to conceptual user modeling. ACM Transactions on Accessible Computing, 15(4):34:1–34:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3551646.

Staff:2011:R

[oACS11] ACM Transactions on Accessible Computing Staff. Reviewers. ACM Transactions on Accessible Computing, 4(1):5:1–5:??, November 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Oh:2015:ABF

[OBFK15] Uran Oh, Stacy Branham, Leah Findlater, and Shaun K. Kane. Audio-based feedback techniques for teaching touchscreen gestures. ACM Transactions on Accessible Computing, 7(3):9:1–9:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Oh:2015:PCH

[OF15] Uran Oh and Leah Findlater. A performance comparison of on-hand versus on-phone nonvisual input by blind and sighted users. ACM Transactions on Accessible Computing, 7(4):14:1– 14:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Obrien:2012:UAT

[ORF12] Marita A. O’Brien, Wendy A. Rogers, and Arthur D. Fisk. Understanding age and technology experience differences in use of prior knowledge for everyday technology interactions. ACM Transactions on Accessible Computing, 4(2):9:1–9:??, March 2012. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Presti:2021:IDS

[PAD^{+}21] Giorgio Presti, Dragan Ahmetovic, Mattia Ducci, Cristian Bernareggi, Luca A. Ludovico, Adriano Baratè, Federico Avanzini, and Sergio Mascetti. Iterative design of sonification techniques to support people with visual impairments in obstacle avoidance. ACM Transactions on Accessible Computing, 14(4):19:1–19:27, December 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3470649.

Prescher:2017:CTP

[PBW17] Denise Prescher, Jens Bornschein, and Gerhard Weber. Consistency of a tactile pattern set. ACM Transactions on Accessible Computing, 10(2):7:1–7:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Portet:2015:PSL

[PCRA15] Franç Ois Portet, Heidi Christensen, Frank Rudzicz, and Jan Alexandersson. Perspectives on speech and language interaction for daily assistive technology: Overall introduction to the special IssuePart 3. ACM Transactions on Accessible Computing, 7 (2):4:1–4:??, July 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Putnam:2016:BPT

[PDR^{+}16] Cynthia Putnam, Maria Dahman, Emma Rose, Jinghui Cheng, and Glenn Bradford. Best practices for teaching accessibility in university classrooms: Cultivating awareness, understanding, and appreciation for diverse users. ACM Transactions on Accessible Computing, 8(4):13:1–13:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Pedrosa:2015:FDC

[PDWT15] Diogo Pedrosa, Maria Da Graça Pimentel, Amy Wright, and Khai N. Truong. Filteryedping: Design challenges and user performance of dwell-free eye typing. ACM Transactions on Accessible Computing, 6(1):3:1–3:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Pellegrini:2015:AAS

[PFM^{+}15] Thomas Pellegrini, Lionel Fontan, Julie Mauclair, Jérôme Farinas, Charlotte Alazard-Guiu, Marina Robert, and Peggy Gatignol. Automatic assessment of speech capability loss in disordered speech. ACM Transactions on Accessible Computing, 6(3):8:1–8:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Palani:2017:PDL

[PG17] Hari Prasath Palani and Nicholas A. Giudice. Principles for designing large-format refreshable haptic graphics using touchscreen devices: an evaluation of nonvisual panning methods. ACM Transactions on Accessible Computing, 9(3):9:1–9:??, February 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Philips:2023:HHI

[PHB23] Gavin R. Philips, Morris Huang, and Cathy Bodine. Helping or hindering: Inclusive design of automated task prompting for workers with cognitive disabilities. ACM Transactions on Accessible Computing, 16(4):24:1–24:??, December 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3628447.

Putnam:2019:ITD

[PHT^{+}19] Cynthia Putnam, Christina Hanschke, Jennifer Todd, Jonathan Gemmell, and Mia Kollia. Interactive technologies designed for children with autism: Reports of use and desires from parents, teachers, and therapists. ACM Transactions on Accessible Computing, 12(3):12:1–12:??, September 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3342285.

Peters:2014:ATA

[PHWH14] Christian Peters, Thomas Hermann, Sven Wachsmuth, and Jesse Hoey. Automatic task assistance for people with cognitive disabilities in brushing teeth — a user study with the TEBRA system. ACM Transactions on Accessible Computing, 5(4):10:1– 10:??, March 2014. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Pino:2010:IOS

[PK10] Alexandros Pino and Georgios Kouroupetroglou. ITHACA: An open source framework for building component-based augmentative and alternative communication applications. ACM Transactions on Accessible Computing, 2(4):14:1–14:??, June 2010. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Pandey:2023:ARH

[PKGG23] Vineet Pandey, Nergis C. Khan, Anoopum S. Gupta, and Krzysztof Z. Gajos. Accuracy and reliability of at-home quantification of motor impairments using a computer-based pointing task with children with ataxia-telangiectasia. ACM Transactions on Accessible Computing, 16(1):10:1–10:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3581790.

Parnandi:2015:DRT

[PKL^{+}15] Avinash Parnandi, Virendra Karappa, Tian Lan, Mostafa Shahin, Jacqueline McKechnie, Kirrie Ballard, Beena Ahmed, and Ricardo Gutierrez-Osuna. Development of a remote therapy tool for childhood apraxia of speech. ACM Transactions on Accessible Computing, 7(3):10:1–10:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Pruszynska:2022:TET

[PMJB^{+}22] Magdalena Pruszyńska, Marta Milewska-Jedrzejczak, Igor Bednarski, Piotr Szpakowski, Andrzej Głabiński, and Sławomir Konrad Tadeja. Towards effective telerehabilitation: Assessing effects of applying augmented reality in remote rehabilitation of patients suffering from multiple sclerosis. ACM Transactions on Accessible Computing, 15(4):37:1–37:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3560822.

Putnam:2023:ICB

[PRM23] Cynthia Putnam, Emma J. Rose, and Craig M. MacDonald. “It could be better. It could be much worse”: Understanding accessibility in user experience practice with implications for industry and education. ACM Transactions on Accessible Computing, 16(1):9:1–9:??, March 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3575662.

Pradhan:2022:DET

[PRR^{+}22] Debashish Pradhan, Tripti Rajput, Aravind Jembu Rajkumar, Jonathan Lazar, Rajiv Jain, Vlad I. Morariu, and Varun Manjunatha. Development and evaluation of a tool for assisting content creators in making PDF files more accessible. ACM Transactions on Accessible Computing, 15(1):3:1–3:52, March 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507661.

Price:2009:DEP

[PS09] Kathleen J. Price and Andrew Sears. The development and evaluation of performance-based functional assessment: a methodology for the measurement of physical capabilities. ACM Transactions on Accessible Computing, 2(2):10:1–10:??, June 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Prietch:2022:SRU

[PSG22] Soraia Prietch, J. Alfredo Sánchez, and Josefina Guerrero. A systematic review of user studies as a basis for the design of systems for automatic sign language processing. ACM Transactions on Accessible Computing, 15(4):36:1–36:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3563395.

Phutane:2022:TMP

[PWC^{+}22] Mahika Phutane, Julie Wright, Brenda Veronica Castro, Lei Shi, Simone R. Stern, Holly M. Lawson, and Shiri Azenkot. Tactile materials in practice: Understanding the experiences of teachers of the visually impaired. ACM Transactions on Accessible Computing, 15(3):17:1–17:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3508364.

Piper:2011:WSA

[PWH11] Anne Marie Piper, Nadir Weibel, and James D. Hollan. Write-N-Speak: Authoring multimodal digital-paper materials for speechlanguage therapy. ACM Transactions on Accessible Computing, 4(1):2:1–2:??, November 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Regimbal:2024:IOS

[RBKC24] Juliette Regimbal, Jeffrey R. Blum, Cyan Kuo, and Jeremy R. Cooperstock. IMAGE: an open-source, extensible framework for deploying accessible audio and haptic renderings of Web graphics. ACM Transactions on Accessible Computing, 17(2):11:1–11:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3665223.

Rello:2016:EFT

[RBY16] Luz Rello and Ricardo Baeza-Yates. The effect of font type on screen readability by people with dyslexia. ACM Transactions on Accessible Computing, 8(4):15:1–15:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Rudzicz:2015:PSL

[RCPA15] Frank Rudzicz, Heidi Christensen, François Portet, and Jan Alexandersson. Perspectives on speech and language interaction for daily assistive technology: Introduction to Part 2 — speaking and reading aids. ACM Transactions on Accessible Computing, 6 (4):11:1–11:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Reichinger:2018:PYM

[RCW^{+}18] Andreas Reichinger, Helena Garcia Carrizosa, Joanna Wood, Svenja Schröder, Christian Löw, Laura Rosalia Luidolt, Maria Schimkowitsch, Anton Fuhrmann, Stefan Maierhofer, and Werner Purgathofer. Pictures in your mind: Using interactive gesturecontrolled reliefs to explore art. ACM Transactions on Accessible Computing, 11(1):2:1–2:??, April 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Rauschenberger:2024:ISI

[RE24] Maria Rauschenberger and Sukru Eraslan. Introduction to the special issue on W4A’22. ACM Transactions on Accessible Computing, 17(2):9:1–9:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3676149.

Richards:2016:IASa

[Ric16a] John Richards. Introduction to the ASSETS ’14 special issue. ACM Transactions on Accessible Computing, 8(1):1:1–1:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Richards:2016:IASb

[Ric16b] John Richards. Introduction to the ASSETS’14 special issue, Part 2. ACM Transactions on Accessible Computing, 8(3):8:1–8:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Randolph:2010:AFN

[RJ10] Adriane B. Randolph and Melody M. Moore Jackson. Assessing fit of nontraditional assistive technologies. ACM Transactions on Accessible Computing, 2(4):16:1–16:??, June 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Rozado:2017:FHC

[RNL17] David Rozado, Jason Niu, and Martin Lochner. Fast humancomputer interaction by combining gaze pointing and face gestures. ACM Transactions on Accessible Computing, 10(3):10:1– 10:??, August 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Rector:2017:DRW

[RVL^{+}17] Kyle Rector, Roger Vilardaga, Leo Lansky, Kellie Lu, Cynthia L. Bennett, Richard E. Ladner, and Julie A. Kientz. Design and real-world evaluation of eyes-free yoga: an exergame for blind and low-vision exercise. ACM Transactions on Accessible Computing, 9 (4):12:1–12:??, April 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Rudzicz:2015:SIP

[RWBM15] Frank Rudzicz, Rosalie Wang, Momotaz Begum, and Alex Mihailidis. Speech interaction with personal assistive robots supporting aging at home for individuals with Alzheimer’s disease. ACM Transactions on Accessible Computing, 7(2):6:1–6:??, July 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Ross:2020:EIL

[RZFW20] Anne Spencer Ross, Xiaoyi Zhang, James Fogarty, and Jacob O. Wobbrock. An epidemiology-inspired large-scale analysis of Android app accessibility. ACM Transactions on Accessible Computing, 13(1):4:1–4:36, April 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3348797.

Sharma:2020:STA

[SAL^{+}20] Sumita Sharma, Tero Avellan, Juhani Linna, Krishnaveni Achary, Markku Turunen, Jaakko Hakulinen, and Blessin Varkey. Sociotechnical aspirations for children with special needs: a study in two locations — India and Finland. ACM Transactions on Accessible Computing, 13(3):13:1–13:27, August 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3396076.

Shinohara:2018:TSA

[SBPW18] Kristen Shinohara, Cynthia L. Bennett, Wanda Pratt, and Jacob O. Wobbrock. Tenets for social accessibility: Towards humanizing disabled people in design. ACM Transactions on Accessible Computing, 11(1):6:1–6:??, April 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Samanta:2020:VTE

[SC20] Debasis Samanta and Tuhin Chakraborty. VectorEntry: Text entry mechanism using handheld touch-enabled mobile devices for people with visual impairments. ACM Transactions on Accessible Computing, 13(3):12:1–12:29, August 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3406537.

Stearns:2016:EHA

[SDO^{+}16] Lee Stearns, Ruofei Du, Uran Oh, Catherine Jou, Leah Findlater, David A. Ross, and Jon E. Froehlich. Evaluating haptic and auditory directional guidance to assist blind people in reading printed text using finger-mounted cameras. ACM Transactions on Accessible Computing, 9(1):1:1–1:??, November 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Sengpiel:2016:TDH

[Sen16] Michael Sengpiel. Teach or design? How older adults’ use of ticket vending machines could be more effective. ACM Transactions on Accessible Computing, 9(1):2:1–2:??, November 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Savage:2022:GCE

[SFSR^{+}22] Saiph Savage, Claudia Flores-Saviaga, Rachel Rodney, Liliana Savage, Jon Schull, and Jennifer Mankoff. The global care ecosystems of 3D printed assistive devices. ACM Transactions on Accessible Computing, 15(4):31:1–31:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3537676.

SassakiResendeSilva:2024:PNA

[SFW^{+}24] Jorge Sassaki Resende Silva, Paula Christina Figueira Cardoso, Raphael Winckler De Bettio, Daniela Cardoso Tavares, Carlos Alberto Silva, Willian Massami Watanabe, and André Pimenta Freire. In-page navigation aids for screen-reader users with automatic topicalisation and labelling. ACM Transactions on Accessible Computing, 17(2):12:1–12:??, June 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3649223.

Sosa-Garcia:2017:HVR

[SGO17] Joan Sosa-García and Francesca Odone. “Hands on” visual recognition for visually impaired users. ACM Transactions on Accessible Computing, 10(3):8:1–8:??, August 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Shipman:2014:ISL

[SGOM14] Frank M. Shipman, Ricardo Gutierrez-Osuna, and Caio D. D. Monteiro. Identifying sign language videos in video sharing sites. ACM Transactions on Accessible Computing, 5(4):9:1–9:??, March 2014. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Sears:2008:I

[SH08] Andrew Sears and Vicki Hanson. Introduction. ACM Transactions on Accessible Computing, 1(1):1:1–1:??, May 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Sears:2012:RUA

[SH12] Andrew Sears and Vicki L. Hanson. Representing users in accessibility research. ACM Transactions on Accessible Computing, 4 (2):7:1–7:??, March 2012. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Sears:2013:E

[SH13] Andrew Sears and Vicki Hanson. Editorial. ACM Transactions on Accessible Computing, 5(2):3:1–3:??, October 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Shinohara:2021:ISI

[SH21] Kristen Shinohara and Foad Hamidi. Introduction to the special issue on ASSETS’19. ACM Transactions on Accessible Computing, 14(4):18e:1–18e:2, December 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3486212.

Sharit:2011:HPS

[SHN^{+}11] Joseph Sharit, Mario A. Hernandez, Sankaran N. Nair, Thomas Kuhn, and Sara J. Czaja. Health problem solving by older persons using a complex government Web site: Analysis and implications for Web design. ACM Transactions on Accessible Computing, 3 (3):11:1–11:??, April 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Shinohara:2020:DSA

[SJPW20] Kristen Shinohara, Nayeri Jacobo, Wanda Pratt, and Jacob O. Wobbrock. Design for social accessibility method cards: Engaging users and reflecting on social scenarios for accessible design. ACM Transactions on Accessible Computing, 12(4):1–33, January 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/3369903.

Sauer:2010:TUU

[SLHF10] Graig Sauer, Jonathan Lazar, Harry Hochheiser, and Jinjuan Feng. Towards a universally usable human interaction proof: Evaluation of task completion strategies. ACM Transactions on Accessible Computing, 2(4):15:1–15:??, June 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Sultana:2019:EAS

[SM19] Afroza Sultana and Karyn Moffatt. Effects of aging on small target selection with touch input. ACM Transactions on Accessible Computing, 12(1):1:1–1:??, February 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3300178.

Sharit:2019:EEU

[SMB^{+}19] Joseph Sharit, Jerad H. Moxley, Walter R. Boot, Neil Charness, Wendy A. Rogers, and Sara J. Czaja. Effects of extended use of an age-friendly computer system on assessments of computer proficiency, attitudes, and usability by older non–computer users. ACM Transactions on Accessible Computing, 12(2):9:1–9:??, July 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

URL https://dl.acm.org/ft_gateway.cfm?id=3325290.

Sharma:2022:CTS

[SMB22] Vinay Krishna Sharma, L. R. D. Murthy, and Pradipta Biswas. Comparing two safe distance maintenance algorithms for a gazecontrolled HRI involving users with SSMI. ACM Transactions on Accessible Computing, 15(3):27:1–27:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3530822.

Sato:2019:NWL

[SOG^{+}19] Daisuke Sato, Uran Oh, João Guerreiro, Dragan Ahmetovic, Kakuya Naito, Hironobu Takagi, Kris M. Kitani, and Chieko Asakawa. NavCog3 in the wild: Large-scale blind indoor navigation assistant with semantic features. ACM Transactions on Accessible Computing, 12(3):14:1–14:??, September 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/ft_gateway.cfm?id=3340319.

Saggion:2015:MIS

[SSB^{+}15] Horacio Saggion, Sanja Stajner, Stefan Bott, Simon Mille, Luz Rello, and Biljana Drndarevic. Making it simplext: Implementation and evaluation of a text simplification system for Spanish. ACM Transactions on Accessible Computing, 6(4):14:1–14:??, June 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Stangl:2022:PCV

[SSD^{+}22] Abigale Stangl, Kristina Shiroma, Nathan Davis, Bo Xie, Kenneth R. Fleischmann, Leah Findlater, and Danna Gurari. Privacy concerns for visual assistance technologies. ACM Transactions on Accessible Computing, 15(2):15:1–15:43, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3517384.

Sanchez:2010:UMV

[SSG10] Jaime Sánchez, Mauricio Saenz, and Jose Miguel Garrido. Usability of a multimodal video game to improve navigation skills for blind children. ACM Transactions on Accessible Computing, 3(2): 7:1–7:??, November 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Snider:2020:AIN

[SST20] Sharon Snider, Willie L. Scott II, and Shari Trewin. Accessibility information needs in the enterprise. ACM Transactions on Accessible Computing, 12(4):1–23, January 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3368620.

Sato:2010:EAC

[STK^{+}10] Daisuke Sato, Hironobu Takagi, Masatomo Kobayashi, Shinya Kawanaka, and Chieko Asakawa. Exploratory analysis of collaborative Web accessibility improvement. ACM Transactions on Accessible Computing, 3(2):5:1–5:??, November 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Sharma:2018:DGB

[SVA^{+}18] Sumita Sharma, Blessin Varkey, Krishnaveni Achary, Jaakko Hakulinen, Markku Turunen, Tomi Heimonen, Saurabh Srivastava, and Nitendra Rajput. Designing gesture-based applications for individuals with developmental disabilities: Guidelines from user studies in India. ACM Transactions on Accessible Computing, 11(1):3:1–3:??, April 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Struve:2009:VMT

[SW09] Doreen Struve and Hartmut Wandke. Video modeling for training older adults to use new technologies. ACM Transactions on Accessible Computing, 2(1):4:1–4:??, May 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Shinohara:2016:SCS

[SW16] Kristen Shinohara and Jacob O. Wobbrock. Self-conscious or selfconfident? A diary study conceptualizing the social accessibility of assistive technology. ACM Transactions on Accessible Computing, 8(2):5:1–5:??, January 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Thevin:2020:XRV

[TBB20] Lauren Thevin, Carine Briant, and Anke M. Brock. X-Road: Virtual reality glasses for orientation and mobility training of people with visual impairments. ACM Transactions on Accessible Computing, 13(2):7:1–7:47, June 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

abs/10.1145/3377879.

Tomlinson:2016:EAG

[TBC^{+}16] Brianna J. Tomlinson, Jared Batterman, Yee Chieh Chew, Ashley Henry, and Bruce N. Walker. Exploring auditory graphing software in the classroom: The effect of auditory graphs on the classroom environment. ACM Transactions on Accessible Computing, 9(1):3:1–3:??, November 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Tartaro:2015:APS

[TCR^{+}15] Andrea Tartaro, Justine Cassell, Corina Ratz, Jennifer Lira, and Valeria Nanclares-Nogués. Accessing peer social interaction: Using authorable virtual peer technology as a component of a group social skills intervention program. ACM Transactions on Accessible Computing, 6(1):2:1–2:??, March 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Tigwell:2017:ACP

[TFA17] Garreth W. Tigwell, David R. Flatla, and Neil D. Archibald. ACE: a colour palette design tool for balancing aesthetics and accessibility. ACM Transactions on Accessible Computing, 9(2):5:1–5:??, January 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Taheri:2024:MET

[TGMBS24] Atieh Taheri, Carlos Gilberto Gomez-Monroy, Vicente Borja, and Misha Sra. MouseClicker: Exploring tactile feedback and physical agency for people with hand motor impairments. ACM Transactions on Accessible Computing, 17(1):5:1–5:??, March 2024. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3648685.

Trewin:2009:EVM

[TLHC09] Shari Trewin, Mark Laff, Vicki Hanson, and Anna Cavender. Exploring visual and motor accessibility in navigating a virtual world. ACM Transactions on Accessible Computing, 2(2):11:1– 11:??, June 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Trinh:2023:EEM

[TMG23] Viet Trinh, Roberto Manduchi, and Nicholas A. Giudice. Experimental evaluation of multi-scale tactile maps created with SIM, a Web app for indoor map authoring. ACM Transactions on Accessible Computing, 16(2):13:1–13:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3590775.

Trnka:2009:UIW

[TMY^{+}09] Keith Trnka, John McCaw, Debra Yarrington, Kathleen F. Mc-Coy, and Christopher Pennington. User interaction with word prediction: The effects of prediction quality. ACM Transactions on Accessible Computing, 1(3):17:1–17:??, February 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Tamburro:2022:CSC

[TNRW22] Carla Tamburro, Timothy Neate, Abi Roper, and Stephanie Wilson. Comic Spin: a comic creation tool enabling self-expression for people with aphasia. ACM Transactions on Accessible Computing, 15(2):8:1–8:27, June 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3508500.

Trewin:2008:GE

[Tre08] Shari Trewin. Guest editorial. ACM Transactions on Accessible Computing, 1(1):2:1–2:??, May 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Tran:2015:EIB

[TRLW15] Jessica J. Tran, Eve A. Riskin, Richard E. Ladner, and Jacob O. Wobbrock. Evaluating intelligibility and battery drain of mobile sign language video transmitted at low frame rates and bit rates. ACM Transactions on Accessible Computing, 7(3):11:1– 11:??, November 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Tang:2024:ESP

[TST^{+}24] Charlotte Tang, Imrul K. Shuva, Matthew Thelen, Linda Zhu, and Nathaniel S. Miller. Exploring the strategies people with Parkinson’s disease use to self-track symptoms and medications. ACM Transactions on Accessible Computing, 17(1):3:1–3:??, March 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3649454.

Vanderheiden:2008:UAC

[Van08] Gregg C. Vanderheiden. Ubiquitous accessibility, common technology core, and micro-assistive technology: Commentary on “Computers and People with Disabilities”. ACM Transactions on Accessible Computing, 1(2):10:1–10:??, October 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Vacher:2019:MEC

[VAR^{+}19] Michel Vacher, Frédéric Aman, Solange Rossato, François Portet, and Benjamin Lecouteux. Making emergency calls more accessible to older adults through a hands-free speech interface in the house. ACM Transactions on Accessible Computing, 12(2):8:1– 8:??, July 2019. CODEN ???? ISSN 1936-7228 (print), 1936-

7236 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3310132.

Vacher:2015:ECA

[VCP^{+}15] Michel Vacher, Sybille Caffiau, François Portet, Brigitte Meillon, Camille Roux, Elena Elias, Benjamin Lecouteux, and Pedro Chahuara. Evaluation of a context-aware voice interface for ambient assisted living: Qualitative user study vs. quantitative system evaluation. ACM Transactions on Accessible Computing, 7(2):5:1– 5:??, July 2015. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

vanErp:2020:TWM

[vEPM20] Jan B. F. van Erp, Katja I. Paul, and Tina Mioch. Tactile working memory capacity of users who are blind in an electronic travel aid application with a vibration belt. ACM Transactions on Accessible Computing, 13(2):5:1–5:14, June 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3372273.

Vargemidis:2021:IGS

[VGV^{+}21] Dimitri Vargemidis, Kathrin Gerling, Vero Vanden Abeele, Luc Geurts, and Katta Spiel. Irrelevant gadgets or a source of worry: Exploring wearable activity trackers with older adults. ACM Transactions on Accessible Computing, 14(3):16:1–16:28, September 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3473463.

Vargemidis:2023:PPE

[VGVG23] Dimitri Vargemidis, Kathrin Gerling, Vero Vanden Abeele, and Luc Geurts. Performance and pleasure: Exploring the perceived usefulness and appeal of physical activity data visualizations with older adults. ACM Transactions on Accessible Computing, 16 (3):21:1–21:??, September 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3615664.

Vickers:2013:PLT

[VIH13] Stephen Vickers, Howell Istance, and Aulikki Hyrskykari. Performing locomotion tasks in immersive computer games with an adapted eye-tracking interface. ACM Transactions on Accessible Computing, 5(1):2:1–2:??, September 2013. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Venkatasubramanian:2023:DPA

[VRL^{+}23] Krishna Venkatasubramanian, Tina-Marie Ranalli, Jack Lanoie, Alexander Sinapi, Andrew Laraw Lama, Jeanine Skorinko, Mariah Freark, and Nancy Alterio. The design and prototyping of an app to teach adults with intellectual and developmental disabilities to empower them against abuse. ACM Transactions on Accessible Computing, 16(2):16:1–16:??, June 2023. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3569585.

VandenAbeele:2021:IVR

[VSH^{+}21] Vero Vanden Abeele, Brenda Schraepen, Hanne Huygelier, Celine Gillebert, Kathrin Gerling, and Raymond Van Ee. Immersive virtual reality for older adults: Empirically grounded design guidelines. ACM Transactions on Accessible Computing, 14(3): 14:1–14:30, September 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3470743.

Williams:2022:MME

[WAMG22] Rua M. Williams, Kiana Alikhademi, Imani N. S. Munyaka, and Juan E. Gilbert. MetaCogs: Mitigating executive dysfunction via agent-based modeling for metacognitive strategy development. ACM Transactions on Accessible Computing, 15(3):24:1–24:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3514254.

Wandmacher:2008:SAC

[WAPD08] Tonio Wandmacher, Jean-Yves Antoine, Franck Poirier, and Jean-Paul Départe. Sibylle, an assistive communication system adapting to the context and its user. ACM Transactions on Accessible Computing, 1(1):6:1–6:??, May 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Waller:2009:ESP

[WBO^{+}09] Annalu Waller, Rolf Black, David A. O’Mara, Helen Pain, Graeme Ritchie, and Ruli Manurung. Evaluating the STANDUP pun generating software with children with cerebral palsy. ACM Transactions on Accessible Computing, 1(3):16:1–16:??, February 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Williams:2019:FSA

[WCG^{+}19] Kristin Williams, Taylor Clarke, Steve Gardiner, John Zimmerman, and Anthony Tomasic. Find and seek: Assessing the impact of table navigation on information look-up with a screen reader. ACM Transactions on Accessible Computing, 12(3):11:1– 11:??, September 2019. CODEN ???? ISSN 1936-7228 (print),

1936-7236 (electronic). URL https://dl.acm.org/ft_gateway.
cfm?id=3342282.

Wang:2016:ISL

[WCH^{+}16] Hanjie Wang, Xiujuan Chai, Xiaopeng Hong, Guoying Zhao, and Xilin Chen. Isolated sign language recognition with Grassmann covariance matrices. ACM Transactions on Accessible Computing, 8(4):14:1–14:??, May 2016. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Wood:2021:IBP

[WDEA^{+}21] Rachel Wood, Emma Dixon, Salma Elsayed-Ali, Ekta Shokeen, Amanda Lazar, and Jonathan Lazar. Investigating best practices for remote summative usability testing with people with mild to moderate dementia. ACM Transactions on Accessible Computing, 14(3):11:1–11:26, September 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3460942.

Wood:2024:HDV

[WFL24] Rachel Wood, Jinjuan Heidi Feng, and Jonathan Lazar. Health data visualization literacy skills of Young adults with down syndrome and the barriers to inference-making. ACM Transactions on Accessible Computing, 17(1):4:1–4:??, March 2024. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3648621.

Wobbrock:2008:GCM

[WG08] Jacob O. Wobbrock and Krzysztof Z. Gajos. Goal crossing with mice and trackballs for people with motor impairments: Performance, submovements, and design directions. ACM Transactions on Accessible Computing, 1(1):4:1–4:??, May 2008. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Wolters:2009:BOD

[WGMM09] Maria Wolters, Kallirroi Georgila, Johanna D. Moore, and Sarah E. MacPherson. Being old doesn’t mean acting old: How older users interact with spoken dialog systems. ACM Transactions on Accessible Computing, 2(1):2:1–2:??, May 2009. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Webster:2014:TSC

[WH14] Gemma Webster and Vicki L. Hanson. Technology for supporting care staff in residential homes. ACM Transactions on Accessible Computing, 5(3):8:1–8:??, January 2014. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Wobbrock:2011:ABD

[WKG^{+}11] Jacob O. Wobbrock, Shaun K. Kane, Krzysztof Z. Gajos, Susumu Harada, and Jon Froehlich. Ability-based design: Concept, principles and examples. ACM Transactions on Accessible Computing, 3(3):9:1–9:??, April 2011. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Walker:2010:UDA

[WM10] B. N. Walker and L. M. Mauney. Universal design of auditory graphs: a comparison of sonification mappings for visually impaired and sighted listeners. ACM Transactions on Accessible Computing, 2(3):12:1–12:??, March 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Wade:2018:EEI

[WNI^{+}18] Joshua Wade, Heathman S. Nichols, Megan Ichinose, Dayi Bian, Esube Bekele, Matthew Snodgress, Ashwaq Zaini Amat, Eric Granholm, Sohee Park, and Nilanjan Sarkar. Extraction of emotional information via visual scanning patterns: a feasibility study of participants with schizophrenia and neurotypical individuals. ACM Transactions on Accessible Computing, 11(4):23:1–23:??, November 2018. CODEN ???? ISSN 1936-7228 (print), 1936-

7236 (electronic). URL https://dl.acm.org/ft_gateway.cfm?
id=3282434.

Wade:2017:PMD

[WSS^{+}17] Joshua Wade, Arpan Sarkar, Amy Swanson, Amy Weitlauf, Zachary Warren, and Nilanjan Sarkar. Process measures of dyadic collaborative interaction for social skills intervention in individuals with autism spectrum disorders. ACM Transactions on Accessible Computing, 10(4):13:1–13:??, October 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Yesilada:2012:SIA

[Yes12] Yeliz Yesilada. Special issue ASSETS 2011. ACM Transactions on Accessible Computing, 4(3):12:1–12:??, December 2012. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Yamagami:2022:TOD

[YJGF^{+}22] Momona Yamagami, Sasa Junuzovic, Mar Gonzalez-Franco, Eyal Ofek, Edward Cutrell, John R. Porter, Andrew D. Wilson, and Martez E. Mott. Two-in-one: a design space for mapping unimanual input into bimanual interactions in VR for users with limited movement. ACM Transactions on Accessible Computing, 15 (3):23:1–23:??, September 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/doi/

10.1145/3510463.

Yamagami:2022:IJO

[YMMS22] Momona Yamagami, Kelly Mack, Jennifer Mankoff, and Katherine M. Steele. “I’m Just Overwhelmed”: Investigating physical therapy accessibility and technology interventions for people with disabilities and/or chronic conditions. ACM Transactions on Accessible Computing, 15(4):35:1–35:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https:

//dl.acm.org/doi/10.1145/3563396.

Yan:2019:CSA

[YR19] Shunguo Yan and P. G. Ramachandran. The current status of accessibility in mobile apps. ACM Transactions on Accessible Computing, 12(1):3:1–3:??, February 2019. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/ft_gateway.cfm?id=3300176.

Zhao:2021:CHC

[ZAM^{+}21] Huan Zhao, Ashwaq Zaini Amat, Miroslava Migovich, Amy Swanson, Amy S. Weitlauf, Zachary Warren, and Nilanjan Sarkar. C-Hg: a collaborative haptic-gripper fine motor skill training system for children with autism spectrum disorder. ACM Transactions on Accessible Computing, 14(2):9:1–9:28, July 2021. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://

dl.acm.org/doi/10.1145/3459608.

Zhao:2022:IHI

[ZAM^{+}22] Huan Zhao, Ashwaq Zaini Amat, Miroslava Migovich, Amy Swanson, Amy S. Weitlauf, Zachary Warren, and Nilanjan Sarkar. INC-Hg: an intelligent collaborative haptic-gripper virtual reality system. ACM Transactions on Accessible Computing, 15(1):5:1–5:23, March 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236

(electronic). URL https://dl.acm.org/doi/10.1145/3487606.

Zhang:2017:MPH

[ZDW17] Ting Zhang, Bradley S. Duerstock, and Juan P. Wachs. Multimodal perception of histological images for persons who are blind or visually impaired. ACM Transactions on Accessible Computing, 9(3):7:1–7:??, February 2017. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Zhu:2010:IGB

[ZFS10] Shaojian Zhu, Jinjuan Feng, and Andrew Sears. Investigating grid-based navigation: The impact of physical disability. ACM Transactions on Accessible Computing, 3(1):3:1–3:??, September 2010. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Zhang:2018:DEC

[ZFS^{+}18] Lian Zhang, Qiang Fu, Amy Swanson, Amy Weitlauf, Zachary Warren, and Nilanjan Sarkar. Design and evaluation of a collaborative virtual environment (CoMove) for autism spectrum disorder intervention. ACM Transactions on Accessible Computing, 11(2): 11:1–11:??, June 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).

Zhang:2022:IOL

[ZMN^{+}22] Han Zhang, Margaret Morris, Paula Nurius, Kelly Mack, Jennifer Brown, Kevin Kuehn, Yasaman Sefidgar, Xuhai Xu, Eve Riskin, Anind Dey, and Jennifer Mankoff. Impact of online learning in the context of COVID-19 on undergraduates with disabilities and mental health concerns. ACM Transactions on Accessible Computing, 15(4):29:1–29:??, December 2022. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/10.1145/3538514.

Zheng:2021:CNI

[ZSS^{+}21] Z. Kevin Zheng, Nandan Sarkar, Amy Swanson, Amy Weitlauf, Zachary Warren, and Nilanjan Sarkar. CheerBrush: a novel interactive augmented reality coaching system for toothbrushing skills in children with autism spectrum disorder. ACM Transactions on Accessible Computing, 14(4):23:1–23:20, December 2021. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL

https://dl.acm.org/doi/10.1145/3481642.

Zhao:2020:DEC

[ZSSA20] Yuhang Zhao, Sarit Szpiro, Lei Shi, and Shiri Azenkot. Designing and evaluating a customizable head-mounted vision enhancement system for people with low vision. ACM Transactions on Accessible Computing, 12(4):1–46, January 2020. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3361866.

Zhao:2018:DHG

[ZZS^{+}18] Huan Zhao, Zhaobo Zheng, Amy Swanson, Amy Weitlauf, Zachary Warren, and Nilanjan Sarkar. Design of a haptic-gripper virtual reality system (Hg) for analyzing fine motor behaviors in children with autism. ACM Transactions on Accessible Computing, 11(4):19:1–19:??, November 2018. CODEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic). URL https://dl.acm.org/ft_

gateway.cfm?id=3231938.

Zhang:2017:PAW

[ZZUK17] Dongsong Zhang, Lina Zhou, Judith O. Uchidiuno, and Isil Y. Kilic. Personalized assistive Web for improving mobile Web browsing and accessibility for visually impaired users. ACM Transactions on Accessible Computing, 10(2):6:1–6:??, April 2017. CO-DEN ???? ISSN 1936-7228 (print), 1936-7236 (electronic).